### v1

# Oriented Grassmannian Bundle, Normal Curvature Reduction, and Persistent Homology

Abstract
We consider a smooth closed orientable submanifold $M\subset\mathbb{R}^{D}$ with narrow cycles. We embed $M$ into a scaled oriented Grassmannian bundle via the Gauss map in order to enlarge the scale of these cycles. Under mild assumptions, we show that this embedding reduces the normal curvature of the embedded submanifold in directions where the original normal curvature is large. For smooth closed hypersurfaces, we further show that this construction increases the distance between antipodal points of narrow cycles for fixed volume.
We then obtain an explicit range of radii for which the ambient Čech complex on this Grassmannian bundle is homotopy equivalent to the embedded manifold, yielding lower bounds on the scales at which the Čech filtration recovers the homology of $M$. Since the distance induced by the embedding depends on both positions and oriented tangent spaces, we work with Whitney $C^{1}$ convergence of embeddings and prove that the associated Čech persistent homology is stable with respect to the interleaving distance. Finally, we describe a procedure for computing a distance matrix for a finite subset with respect to this embedding and illustrate the construction on several examples, including an approximate quasi-halo orbit in the Saturn–Enceladus system.
1 Introduction
Persistent homology computes barcodes of filtered chain complexes constructed from a metric space $X$ [zomorodian2004computing, oudot2015persistence]. For $r>0$, the Čech complex $\check{C}(X;r)$ is the nerve of the open cover of $r$-balls $\{B_{X}(x,r)\}_{x\in X}$, and the Vietoris–Rips complex $\mathrm{VR}(X;r)$ is the abstract simplicial complex whose simplices are the finite subsets $\sigma\subset X$ with $\operatorname{diam}(\sigma)<r$. If $M$ is a Riemannian manifold, then the Čech complex $\check{C}(M;r)$ is homotopy equivalent to $M$ for $r<r_{\operatorname{conv}}(M)$ by the nerve lemma, while the systole and the filling radius bound the range of $r$ for which the Vietoris–Rips complex $\mathrm{VR}(M;r)$ has the same first and top-dimensional homology groups as $M$ [virk20201, lim2024vietoris].
The reach (or normal injectivity radius) $\mathsf{rch}_{\mathbb{R}^{D}}(M)$ of a smooth submanifold $M\subset\mathbb{R}^{D}$ is the supremum of $\tau>0$ such that every point $x\in\mathbb{R}^{D}$ with $d_{\mathbb{R}^{D}}(x,M)<\tau$ has a unique nearest point in $M$ [federer1959curvature, aamari2019estimating]. For smooth embedded submanifolds, this agrees with the supremum of radii for which the normal exponential map is injective. The reach of $M$ gives bounds on the scales $r$ where the Vietoris–Rips or Čech complex built from sufficiently dense points of $M$ in the ambient Euclidean metric [niyogi2008finding, kim2019homotopy], and the metric thickening of $M$ [adams2019metric, adams2022metric], the 1-Wasserstein analogue of the Vietoris–Rips complex, are homotopy equivalent to $M$.
Aamari et al. [aamari2019estimating] showed that the reach of a submanifold $M\subset\mathbb{R}^{D}$ can be expressed as the minimum of the reciprocal of the maximal normal curvature of $M$ and the minimal width of its bottlenecks, pairs of points in $M$ whose connecting line segment is orthogonal to both tangent spaces. For example, for a highly eccentric ellipse in $\mathbb{R}^{2}$ or a thin $2$-torus with small systole in $\mathbb{R}^{3}$ as in Figure 1, the normal exponential map ceases to be injective at a radius comparable to the reach, since the points of large curvature (marked in red) induce focal points and the bottlenecks (marked in blue) lie close to each other in the Euclidean metric.
In both examples the indicated bottlenecks have opposite normal directions. This suggests that, for an orientable submanifold $M\subset\mathbb{R}^{D}$, embedding $M$ into an augmented ambient space, such as the oriented Grassmannian bundle $\mathbb{R}^{D}\times\mathbf{Gr}^{+}(D,d)$, can enlarge the normal injectivity radius of $M$. In the product metric on $\mathbb{R}^{D}\times\mathbf{Gr}^{+}(D,d)$, the contribution from the Grassmannian factor enlarges the lengths of curves near points of large normal curvature, so that unit-speed geodesics have smaller normal curvature, and bottleneck pairs with opposite normal directions are pushed farther apart because their oriented tangent spaces are far from each other in $\mathbf{Gr}^{+}(D,d)$.
Figure 1: A highly eccentric ellipse (left) and a thin 2-torus (right). Points of large curvature are marked in red, and the narrowest bottlenecks are indicated in blue.
To formalize the above discussion, let $\iota:M\to\mathbb{R}^{D}$ be an embedding of a smooth closed orientable $d$-dimensional submanifold. Fix $c>0$ and consider the scaled oriented Grassmannian $(\mathbf{Gr}_{c}^{+}(D,d),g^{\mathbf{Gr}_{c}^{+}(D,d)})$, whose underlying manifold is the oriented Grassmannian $\mathbf{Gr}^{+}(D,d)$ equipped with the rescaled metric $g^{\mathbf{Gr}_{c}^{+}(D,d)}=c\cdot g^{\mathbf{Gr}^{+}(D,d)}$. Let $\bar{\mathbf{g}}^{+}$ denote the Gauss map, and define the rescaled Gauss map
$\bar{\mathbf{g}}^{+}_{c}:M\xrightarrow{\bar{\mathbf{g}}^{+}}(\mathbf{Gr}^{+}(D,d),g^{\mathbf{Gr}^{+}(D,d)})\xrightarrow{\cdot c}(\mathbf{Gr}_{c}^{+}(D,d),g^{\mathbf{Gr}_{c}^{+}(D,d)}).$
We consider the graph embedding $\iota\times\bar{\mathbf{g}}_{c}^{+}:M\to(\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d),\,g^{\mathbb{R}^{D}}\oplus g^{\mathbf{Gr}_{c}^{+}(D,d)}),$ and define $g_{c}=(\iota\times\bar{\mathbf{g}}^{+}_{c})^{*}(g^{\mathbb{R}^{D}}\oplus g^{\mathbf{Gr}_{c}^{+}(D,d)})$. Throughout, we write $\mathbf{II}$ for the second fundamental form of $M\subset\mathbb{R}^{D}$ with the induced Euclidean metric and $\mathbf{II}_{c}$ for that of $(M,g_{c})\subset\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$ with the product metric. Likewise, we write $\mathrm{vol}_{c}(M)$ for the Riemannian volume of $(M,g_{c})$ and $\mathrm{vol}(M)$ for the Euclidean volume of $M\subset\mathbb{R}^{D}$.
Definition 1.1.
For $c>0$, we define a distance $d_{c}$ on $M$ by
$d_{c}(q_{1},q_{2})=d_{\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)}\bigl((q_{1},\bar{\mathbf{g}}_{c}^{+}(q_{1})),(q_{2},\bar{\mathbf{g}}_{c}^{+}(q_{2}))\bigr)=\sqrt{\|q_{1}-q_{2}\|_{\mathbb{R}^{D}}^{2}+c\,d_{\mathbf{Gr}^{+}(D,d)}\bigl(\bar{\mathbf{g}}^{+}(q_{1}),\bar{\mathbf{g}}^{+}(q_{2})\bigr)^{2}},$
where $\|\cdot\|_{\mathbb{R}^{D}}$ denotes the Euclidean norm and $d_{\mathbf{Gr}^{+}(D,d)}$ is the Riemannian distance on the oriented Grassmannian $\mathbf{Gr}^{+}(D,d)$.
We first show that, by embedding $M$ into $\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$, we can reduce the normal curvature of $M$ along directions where the original normal curvature is large. For $p\in M$ and a nonzero vector $v\in T_{p}M$, denote by $\kappa_{c}(v)$ and $\kappa(v)$ the (operator norms of the) normal curvatures of $(M,g_{c})\subset\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$ at $(p,\bar{\mathbf{g}}_{c}^{+}(p))$ in the direction $(v,d\bar{\mathbf{g}}_{c}^{+}(v))$ and of $M\subset\mathbb{R}^{D}$ at $p$ in the direction $v$, respectively, where both geodesics are parametrized by arc length.
Theorem 1.1.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable $d$-dimensional submanifold. Fix $p\in M$ and a unit vector $v\in T_{p}M$. Assume that the normal curvature at $p$ in the direction $v$ is sufficiently large so that
$\mathrm{vol}_{c}(M)\leq\bigl(1+c\,\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}\bigr)^{d/2}\mathrm{vol}(M),\qquad\kappa(v)>\frac{\|\nabla_{v}^{M}\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}}{\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}},$
where $\|\cdot\|_{\operatorname{HS}}$ denotes the Hilbert–Schmidt norm. Then, for every $c>0$,
$\kappa_{c}(v)\,\mathrm{vol}_{c}(M)^{1/d}\;\leq\;\frac{\sqrt{\kappa(v)^{2}+c\|\nabla_{v}^{M}\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}}{1+c\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}\,\mathrm{vol}_{c}(M)^{1/d}\;<\;\kappa(v)\,\mathrm{vol}(M)^{1/d},$
and the middle term in the inequalities is strictly decreasing as $c$ grows.
In directions with large normal curvatures, even a small displacement along the curve already produces a large change in the tangent space, so the Grassmannian component of $g_{c}$ contributes an increasingly large portion of the speed. Since the normal curvature is defined as the size of the normal acceleration divided by the square of this speed, reparameterizing the curve to have unit speed with respect to $g_{c}$ makes the normal curvature decrease as $c$ grows.
Next, to quantify how distances between bottleneck pairs change with respect to $d_{c}$, we set
$L_{c}(M)$ $=\min\Bigl\{\frac{1}{2}d_{c}(q_{1},q_{2})\;\Big|\;(q_{1},q_{2})\text{ is a bottleneck of }M\subset\mathbb{R}^{D}\Bigr\}$
$=\min\Bigl\{\frac{1}{2}d_{c}(q_{1},q_{2})\;\Big|\;\overline{q_{1}q_{2}}\perp T_{q_{1}}M,\ \overline{q_{1}q_{2}}\perp T_{q_{2}}M\Bigr\}.$
Similarly, in the Euclidean case we define
$L(M)=\min\Bigl\{\frac{1}{2}\|q_{1}-q_{2}\|_{\mathbb{R}^{D}}\;\Big|\;(q_{1},q_{2})\text{ is a bottleneck of }M\subset\mathbb{R}^{D}\Bigr\}.$
Under a mild curvature hypothesis, we show that for a smooth closed orientable hypersurface $M$, the bottlenecks are separated when distances are measured with $d_{c}$ from Definition 1.1.
Theorem 1.2.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable hypersurface. Suppose that $L(M)\leq\frac{1}{\|\mathbf{II}\|_{2}}$. Then for every $c\in\bigl(0,\tfrac{12L(M)^{2}}{\pi^{2}}\bigr]$,
$\frac{L_{c}(M)}{\mathrm{vol}_{c}(M)^{1/(D-1)}}\;\geq\;\frac{\sqrt{4L(M)^{2}+c\pi^{2}}}{2\,\mathrm{vol}_{c}(M)^{1/(D-1)}}\;>\;\frac{L(M)}{\mathrm{vol}(M)^{1/(D-1)}},$
and the middle term in the inequalities is strictly increasing in $c$ on the interval $\bigl(0,\tfrac{12L(M)^{2}}{\pi^{2}}\bigr]$.
We obtain an explicit range of radii $r$ for which the ambient Čech complex $\check{C}(M,\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d);r)$ is homotopy equivalent to $M$. In particular, this yields a lower bound on the lengths of the barcodes in the associated persistent homology, from which the homology of $M$ is recovered from the distance $d_{c}$.
Theorem 1.3.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable $d$-dimensional submanifold. For $c>0$, the ambient Čech complex $\check{C}(M,\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d);r)$ is homotopy equivalent to $M$ for all
$r<\min\!\left(\sqrt{\frac{c}{2}}\arctan\sqrt{\frac{2}{c\|\mathbf{II}_{c}\|_{2}}},\,\frac{\sqrt{c}\,\pi}{2},\,L^{\prime}_{c}(M)\right),$
where
$L^{\prime}_{c}(M)=\min\Bigl\{\frac{1}{2}d_{c}(q_{1},q_{2})\;\Bigm|\;\overline{(q_{1},\bar{\mathbf{g}}_{c}^{+}(q_{1}))(q_{2},\bar{\mathbf{g}}_{c}^{+}(q_{2}))}\perp T_{(q_{i},\bar{\mathbf{g}}_{c}^{+}(q_{i}))}(M,g_{c}),\ i=1,2\Bigr\}.$
[…]
Theorem 1.4.
Let $\{M_{i}\}_{i\in\mathbb{N}}$ be a sequence of smooth closed orientable $d$-dimensional submanifolds of $\mathbb{R}^{D}$ converging to a smooth closed orientable submanifold $M_{\infty}\subset\mathbb{R}^{D}$ in the Whitney $C^{1}$ topology (see Definition 2.1). Fix $c>0$. Then for every field $\mathbb{k}$ and every $j\in\mathbb{N}\cup\{0\}$,
[…]
where $d_{I}$ denotes the interleaving distance between persistence modules.
Based on the above theorems, we describe a method for computing the distance matrix for a finite subset of $M$ with respect to the distance $d_{c}$ from Definition 1.1. To illustrate the performance of this method, we display three computational examples: a time-delay embedded attractor, an approximate quasi-halo orbit in the Saturn–Enceladus system, and a classification of three-dimensional image shapes.
2 Theoretical background
2.1 Persistence theory
A filtered space (or filtration) $\mathbb{X}$ is a collection of topological spaces $\{X_{i}\}_{i\in\mathbb{R}_{\geq 0}}$ equipped with inclusion maps $\iota_{i}^{j}:X_{i}\to X_{j}$ for all $i\leq j$. Let $(X,d)$ be a compact metric space and let $\varepsilon\in\mathbb{R}_{\geq 0}$. The Vietoris–Rips complex $\mathrm{VR}(X;\varepsilon)$ is the abstract simplicial complex whose simplices are finite subsets $\sigma\subset X$ such that $d(x,y)<\varepsilon$ for all $x,y\in\sigma$. For a non-decreasing sequence of parameters $\varepsilon_{0}\leq\varepsilon_{1}\leq\varepsilon_{2}\leq\dots$, these complexes form a filtered space
$\mathrm{VR}(X;\varepsilon_{0})\hookrightarrow\mathrm{VR}(X;\varepsilon_{1})\hookrightarrow\mathrm{VR}(X;\varepsilon_{2})\hookrightarrow\cdots$
[…]
For $n\in\mathbb{N}\cup\{0\}$, a field $\mathbb{k}$ and a filtration $\mathbb{X}=\{X_{i}\}_{i\in\mathbb{R}_{\geq 0}}$, the $n$-th persistent homology $H_{n}(\mathbb{X};\mathbb{k})$ of $\mathbb{X}$ is a family of vector spaces $H_{n}(X_{i};\mathbb{k})$ with induced linear maps $H_{n}(\iota_{i}^{j}):H_{n}(X_{i};\mathbb{k})\to H_{n}(X_{j};\mathbb{k})$ for $i\leq j$. If $H_{n}(\mathbb{X};\mathbb{k})$ is pointwise finite-dimensional and decomposes as a direct sum of interval modules, then it corresponds to a multiset of points $(b,d)\subset\mathbb{R}_{\geq 0}^{2}$, called the persistence diagram in degree $n$.
For a field $\mathbb{k}$, a persistence module (over $\mathbb{R}_{\geq 0}$) $\mathbb{V}$ is a family of vector spaces $\{V_{i}\}_{i\in\mathbb{R}_{\geq 0}}$ equipped with linear maps $v_{i}^{j}:V_{i}\to V_{j}$ for every $0\leq i\leq j$, satisfying $v_{j}^{k}\circ v_{i}^{j}=v_{i}^{k}$ whenever $i\leq j\leq k$ and $v_{i}^{i}$ is the identity map on $V_{i}$. A morphism of degree $\varepsilon$ between two persistence modules $\mathbb{V}$ and $\mathbb{W}$ is a family of linear maps $\Phi=\{\phi_{i}:V_{i}\to W_{i+\varepsilon}\}_{i\in\mathbb{R}_{\geq 0}}$ such that for all $i\leq j$,
$w_{i+\varepsilon}^{j+\varepsilon}\circ\phi_{i}\;=\;\phi_{j}\circ v_{i}^{j}.$
For $\varepsilon\geq 0$, the shift $\Sigma^{\varepsilon}$ of $\mathbb{V}$ by $\varepsilon$ is
$(\Sigma^{\varepsilon}\mathbb{V})_{i}=V_{i+\varepsilon},\qquad(\Sigma^{\varepsilon})(v_{i}^{j})=v_{i+\varepsilon}^{j+\varepsilon}:V_{i+\varepsilon}\to V_{j+\varepsilon}.$
Two persistence modules $\mathbb{V}$ and $\mathbb{W}$ are $\varepsilon$-interleaved if there exist morphisms $\Phi=\{\phi_{i}:V_{i}\to W_{i+\varepsilon}\}_{i\in\mathbb{R}_{\geq 0}}$ and $\Psi=\{\psi_{i}:W_{i}\to V_{i+\varepsilon}\}_{i\in\mathbb{R}_{\geq 0}}$ of degree $\varepsilon$ such that
$\psi_{\,i+\varepsilon}\circ\phi_{i}\;=\;v_{i}^{i+2\varepsilon},\qquad\phi_{\,i+\varepsilon}\circ\psi_{i}\;=\;w_{i}^{i+2\varepsilon},$
for every $i\in\mathbb{R}_{\geq 0}$. The interleaving distance between $\mathbb{V}$ and $\mathbb{W}$ is
$d_{I}(\mathbb{V},\mathbb{W})\;=\;\inf\bigl\{\varepsilon\geq 0\,\big|\,\mathbb{V}\text{ and }\mathbb{W}\text{ are }\varepsilon\text{-interleaved}\bigr\}.$
For more details, see e.g. [oudot2015persistence, chazal2021introduction].
2.2 Basic Riemannian geometry
In this subsection, we refer to [do1992riemannian] and [lee2018introduction]. Let $M$ be a smooth compact $n$-dimensional manifold. For a choice of a smooth section $g\in\Gamma(M;\mathrm{Sym}^{2}T^{*}M)$, a pair $(M,g)$ is called a Riemannian manifold if for each $p\in M$, the fiberwise bilinear map $g_{p}:T_{p}M\times T_{p}M\to\mathbb{R}$ is positive definite. The section $g$ is called a (Riemannian) metric on $M$. We call the map
$\nabla:\mathfrak{X}(M)\times\mathfrak{X}(M)\to\mathfrak{X}(M);\quad(X,Y)\mapsto\nabla_{X}Y$
an affine connection if $\nabla_{X}Y$ is linear over $C^{\infty}(M)$ in $X$ and linear over $\mathbb{R}$ in $Y$, satisfying $\nabla_{X}(fY)=f\nabla_{X}Y+(Xf)Y$ for any $f\in C^{\infty}(M)$. The Levi–Civita connection is the unique affine connection $\nabla$ on $(M,g)$ that satisfies $\nabla(g(X,Y))=g(\nabla X,Y)+g(X,\nabla Y)$ and $\nabla_{X}Y-\nabla_{Y}X=[X,Y]$ for any $X,Y\in\mathfrak{X}(M).$
For a Riemannian manifold $(M,g)$ with Levi–Civita connection $\nabla$, the Riemann curvature tensor $R$ is defined by $R(X,Y)Z=\nabla_{X}\nabla_{Y}Z-\nabla_{Y}\nabla_{X}Z-\nabla_{[X,Y]}Z$ for $X,Y,Z\in\mathfrak{X}(M)$. For a tangent $2$-plane $\Sigma_{p}M\subset T_{p}M$ with an orthonormal basis $\{u,v\}$, the sectional curvature of $\Sigma_{p}M$ is defined by $K_{p}(u,v)=g_{p}(R(u,v)v,u).$
2.3 Convergence theory
We refer to [gromov2007metric, hirsch2012differential] in this subsection. Let $(Z,d_{Z})$ be a metric space and let $A,B\subset Z$ be nonempty compact subsets. The Hausdorff distance between $A$ and $B$ is defined by
$d_{H}^{Z}(A,B)=\max\Big\{\,\sup_{a\in A}\inf_{b\in B}d_{Z}(a,b)\;,\;\sup_{b\in B}\inf_{a\in A}d_{Z}(a,b)\,\Big\}.$
The following stability theorem establishes the continuity of persistent homology with respect to the Hausdorff distance.
Theorem 2.1 ([chazal2014persistence, Theorem 5.6]).
Let $A,B$ be compact subsets of a metric space $Z$. Then for every $j\in\mathbb{N}\cup\{0\}$ and every field $\mathbb{k}$,
[…]
Let $(N,g_{N})$ be a fixed Riemannian manifold, and let $\{M_{i}\}_{i\in\mathbb{N}}$ be a family of smooth closed submanifolds of $N$ such that $M_{i}=F_{i}(M)$ for every $i$, where $M$ is a fixed smooth closed manifold and each $F_{i}:M\to N$ is a smooth embedding.
Definition 2.1 ([hirsch2012differential, Chapter 2]).
We say that the family of submanifolds $\{M_{i}\}_{i\in\mathbb{N}}$ converges to $M_{\infty}$ in the (Whitney) $C^{k}$ topology if there exists an embedding
$F_{\infty}:M\to N,\qquad M_{\infty}=F_{\infty}(M),$
such that $F_{i}\to F_{\infty}$ in the Whitney $C^{k}$ topology on $C^{k}(M,N)$. That is, there exist a locally finite atlas $\{(U_{\alpha},\varphi_{\alpha})\}_{\alpha}$ of $M$ and coordinate charts $\{(V_{\alpha},\psi_{\alpha})\}_{\alpha}$ of $N$ with $F_{\infty}(U_{\alpha})\subset V_{\alpha}$ for each $\alpha$ such that, for each $\alpha$ and each $\varepsilon_{\alpha}>0$, there exists $n_{\alpha}\in\mathbb{N}$ with the property that for every $i\geq n_{\alpha}$,
$\max_{|\beta|\leq k}\sup_{x\in\varphi_{\alpha}(U_{\alpha})}\left\|D^{\beta}\big(\psi_{\alpha}\circ F_{i}\circ\varphi_{\alpha}^{-1}\big)(x)-D^{\beta}\big(\psi_{\alpha}\circ F_{\infty}\circ\varphi_{\alpha}^{-1}\big)(x)\right\|_{2}<\varepsilon_{\alpha}.$
2.4 Geometry of an embedded submanifold
Let $M\subset\mathbb{R}^{D}$ be a smooth closed $d$-dimensional submanifold with $d<D$. Two distinct points $q_{1},q_{2}\in M$ form a bottleneck in $\mathbb{R}^{D}$ if the line segment $\overline{q_{1}q_{2}}$ is orthogonal to both tangent spaces $T_{q_{1}}M$ and $T_{q_{2}}M$. The width of such a bottleneck is defined as $\tfrac{1}{2}\|q_{1}-q_{2}\|_{\mathbb{R}^{D}}$. This notion should not be confused with the bottleneck distance between persistence modules or with bottlenecks in graph theory.
Define the (Euclidean) reach of $M\subset\mathbb{R}^{D}$ by
[…]
Theorem 2.2 ([aamari2019estimating, Theorem 3.4], [breiding2024metric, Theorem 7.8]).
Suppose that the reach of a closed submanifold $M\subset\mathbb{R}^{D}$ is $\tau>0$. Then at least one of the following holds.
  • There exist a point $q\in M$ and a unit-speed geodesic $\gamma(t)$ on $M$ with $\gamma(0)=q$ and $|\gamma^{\prime\prime}(0)|=\frac{1}{\tau}$.
  • There exist distinct points $q_{1},q_{2}\in M$ forming a bottleneck of $M\subset\mathbb{R}^{D}$ such that $\|q_{1}-q_{2}\|_{\mathbb{R}^{D}}=2\tau$.
In particular, the reach $\tau$ of $M$ is realized either as the reciprocal of a normal curvature or as the width of a bottleneck pair.
Suppose a $d$-dimensional smooth compact manifold $M$ is embedded in a closed Riemannian manifold $(N,g)$. Denote the Levi–Civita connection of $N$ by $\nabla$ and the normal bundle of $M$ by $\nu$. The second fundamental form of $M$ is defined to be the tensor $\mathbf{II}:\mathfrak{X}(M)\times\mathfrak{X}(M)\to\Gamma(\nu)$ given by $\mathbf{II}(X,Y)=(\nabla_{X}Y)^{\perp}$, where the map $(\cdot)^{\perp}$ denotes the orthogonal projection onto $\nu$. The normal exponential map $\exp_{\nu}:\nu\to N$ is defined by
$\exp_{\nu}(q,v)=\exp_{q}(v),\quad(q,v)\in\nu.$
For $p\in M$ and a nonzero vector $v\in T_{p}M$, the (norm of the) normal curvature $\kappa(v)$ at $p$ in the direction $v$ is
$\kappa(v)=\sup_{\begin{subarray}{c}w\in\nu_{p}\\ \|w\|_{2}=1\end{subarray}}\frac{g_{p}\bigl(\mathbf{II}_{p}(v,v),\,w\bigr)}{\|v\|_{2}^{2}}.$
We set $\kappa(0)=0$. The operator norm (or 2-norm) of $\mathbf{II}_{p}$ is
$\|\mathbf{II}_{p}\|_{2}=\sup_{\begin{subarray}{c}v\in T_{p}M\\ \|v\|_{2}=1\end{subarray}}\kappa(v)=\sup_{\begin{subarray}{c}v\in T_{p}M\\ \|v\|_{2}=1\end{subarray}}\bigl\|\mathbf{II}_{p}(v,v)\bigr\|_{2}.$
Denote $\|\mathbf{II}\|_{2}=\sup_{p\in M}\|\mathbf{II}_{p}\|_{2}$. For two linear operators $A_{p},B_{p}:T_{p}M\to\nu_{p}$ and an orthonormal basis $\{e_{i}\}_{i=1}^{d}$ of $T_{p}M$, the Hilbert–Schmidt inner product between them is
$\langle A_{p},B_{p}\rangle_{\operatorname{HS}}=\sum_{i=1}^{d}g_{p}\bigl(A_{p}(e_{i}),B_{p}(e_{i})\bigr).$
Denote by $\|\cdot\|_{\operatorname{HS}}$ the induced norm.
Definition 2.2 ([prasad2023cut, Definition 2.5], [attali2022tight, Definitions 12 and 13]).
Let $S$ be a closed subset embedded in a closed Riemannian manifold $N$. Denote the distance function on $N$ by $d_{N}$ and the length of a curve $\gamma\subset N$ by $\mathrm{len}_{N}(\gamma)$. We define a geodesic $\gamma:[0,T]\to N$ to be ($S$-)distance-minimal if $\mathrm{len}_{N}(\gamma|_{[0,t]})=d_{N}(S,\gamma(t))$ for all $t\in[0,T]$. The cut locus of $S$, denoted by $\mathsf{Cu}_{N}(S)$, is the set of points $p\in N$ for which there exists a distance-minimal geodesic $\gamma$ from $S$ to $p$ such that any extension of $\gamma$ beyond its endpoint $p$ is not distance-minimal. The (cut locus) reach (or normal injectivity radius) of $S$ in $N$ is
$\mathsf{rch}_{N}(S)=\inf\{d_{N}(p,q)\mid q\in S,\ p\in\mathsf{Cu}_{N}(S)\}.$
If $S$ is a smooth closed submanifold, then $\mathsf{rch}_{N}(S)$ is the supremum of all $\varepsilon>0$ such that the restriction of the normal exponential map
$\exp_{\nu}:\{(p,v)\in\nu\mid\|v\|_{2}<\varepsilon\}\longrightarrow N$
is an embedding. We give an analogue of Theorem 2.2 below.
Theorem 2.3 ([singh1988closest, Section 2], [basu2023connection, Lemma A.2]).
Let $M$ be a smooth closed submanifold of a Riemannian manifold $(N,g)$, and suppose $\mathsf{rch}_{N}(M)=T>0$. Denote the normal bundle of $M$ by $\nu$ and its normal unit bundle by $S(\nu)$. Then one of the following holds (see Figure 2):
  • (focal)
    There exists a pair $(p,v)\in S(\nu)$ such that the differential of the normal exponential map $d(\exp_{\nu})_{(p,Tv)}$ is not of full rank, while $d(\exp_{\nu})_{(p^{\prime},tv^{\prime})}$ has full rank for every $(p^{\prime},v^{\prime})\in S(\nu)$ and $0<t<T$. In this case $\exp_{\nu}(Tv)$ is called a first focal point and $T$ a first focal time.
  • (non-focal)
    There exist at least two distinct unit-speed distance-minimal geodesics $\gamma_{1},\gamma_{2}:[0,T]\to N$ from (not necessarily distinct) two points in $M$ satisfying $\gamma_{1}(T)=\gamma_{2}(T)$ and $\gamma_{i}^{\prime}(0)\perp M$ for $i=1,2$.
Figure 2: Examples of cut loci (in red): a first focal point caused by large normal curvature (left), a non-focal cut locus arising from two distance-minimal geodesics from the same point (middle), and a non-focal cut locus arising from two distinct points (right).
We will establish a lower bound on the reach of the embedding $(M,g_{c})\subset\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$ by deriving casewise lower bounds. It provides a lower bound on the radii $r$ from which the ambient Čech complex $\check{C}(M,\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d);r)$ or the Vietoris–Rips complex $\mathrm{VR}(\mathbf{Y},\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d);r)$ for a finite subset $\mathbf{Y}\subset M$.
Theorem 2.4 ([basu2023connection, Theorem A]).
Denote the function $f=d_{N}(\cdot,M)^{2}:N\to\mathbb{R}$. Then $f$ is a Morse–Bott function on $N\setminus\mathsf{Cu}_{N}(M)$ with critical manifold $M$, and the set $N\setminus\mathsf{Cu}_{N}(M)$ deformation retracts onto $M$ via the gradient flow of $f|_{N\setminus\mathsf{Cu}_{N}(M)}$.
Theorem 2.5 ([attali2022tight, Proposition 16], modified).
Suppose that the ambient manifold $N$ has non-negative sectional curvature. If a finite subset $\{x_{i}\}_{i=1}^{n}\subset S$ is $\varepsilon/2$-dense in $S$ for some $\varepsilon<\sqrt{6-4\sqrt{2}}\;\mathsf{rch}_{N}(S)$, then there exists a nonempty interval of radii $r$ such that the union of balls $\bigcup_{i=1}^{n}B_{N}(x_{i},r)$ deformation retracts onto $S$. Moreover, the length of this interval depends only on $\varepsilon$ and $\mathsf{rch}_{N}(S)$.
2.5 Geometry of the Grassmannian
We mainly follow [bendokat2024grassmann] and [lai2025simple]. The compact manifold $\mathbf{Gr}(D,d)$, called the (real) Grassmannian, is the space of all $d$-dimensional linear subspaces of $\mathbb{R}^{D}$ equipped with the quotient topology. For two elements $P,Q\in\mathbf{Gr}(D,d)$, let $A,B$ be matrices whose columns form orthonormal bases for $P$ and $Q$, respectively. Let $\sigma_{1},\sigma_{2},\dots,\sigma_{d}$ denote the singular values of $A^{\top}B$. The principal angles between $P$ and $Q$ are defined by
$(\theta_{1},\dots,\theta_{d})=(\arccos{\sigma_{1}},\dots,\arccos{\sigma_{d}}).$
The Grassmannian $\mathbf{Gr}(D,d)$ carries a Riemannian metric whose induced geodesic distance is
$d_{\mathbf{Gr}(D,d)}(P,Q)=\left(\sum_{i=1}^{d}\theta^{2}_{i}\right)^{\frac{1}{2}}.$
Analogously, the space of $d$-dimensional oriented subspaces of $\mathbb{R}^{D}$ is called the (real) oriented Grassmannian, denoted by $\mathbf{Gr}^{+}(D,d).$ For two oriented subspaces $P,Q\in\mathbf{Gr}^{+}(D,d)$ with orthonormal basis matrices $A$ and $B$, the geodesic distance between $P$ and $Q$ is given by
$d_{\mathbf{Gr}^{+}(D,d)}(P,Q)=\begin{cases}\left(\displaystyle\sum_{i=1}^{d}\theta_{i}^{2}\right)^{\frac{1}{2}}&\text{if }\det(A^{\top}B)>0,\\[6.00006pt] \left(\displaystyle(\pi-\theta_{1})^{2}+\sum_{i=2}^{d}\theta_{i}^{2}\right)^{\frac{1}{2}}&\text{otherwise},\end{cases}$
where $\theta_{1},\dots,\theta_{d}$ are the principal angles between $P,Q$ with $\theta_{1}\geq\dots\geq\theta_{d}\geq 0$.
For a smooth closed orientable $d$-dimensional submanifold $M\subset\mathbb{R}^{D}$, the (generalized) Gauss map $\bar{\mathbf{g}}^{+}:M\to\mathbf{Gr}^{+}(D,d)$ is defined by the composition
$p\longmapsto T_{p}M\xrightarrow{\;\cong\;}P\in\mathbf{Gr}^{+}(D,d)$
[…]
3 Embedding into the oriented Grassmannian bundle
3.1 Bound for Riemannian volume
Lemma 3.1 ([lai2025simple, Section 3]).
Let $M\subset\mathbb{R}^{D}$ be a smooth closed $d$-dimensional submanifold with normal bundle $\nu$. Denote the unoriented Gauss map by $\bar{\mathbf{g}}:M\to\mathbf{Gr}(D,d)$. For any $X,Y\in\mathfrak{X}(M)$ and $p\in M$,
$d\bar{\mathbf{g}}_{p}(X_{p})=\mathbf{II}_{p}(X_{p},\cdot)\in\operatorname{Hom}(T_{p}M,\nu_{p}),$
where $\mathbf{II}$ denotes the second fundamental form of $M\subset\mathbb{R}^{D}$. Moreover,
$g^{\mathbf{Gr}(D,d)}_{\bar{\mathbf{g}}(p)}\bigl(d\bar{\mathbf{g}}_{p}(X_{p}),d\bar{\mathbf{g}}_{p}(Y_{p})\bigr)=\langle\mathbf{II}_{p}(X_{p},\cdot),\mathbf{II}_{p}(Y_{p},\cdot)\rangle_{\operatorname{HS}}.$
Theorem 3.1.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable $d$-dimensional submanifold, and $\mathbf{II}$ be the second fundamental form of $M$ with $\|\mathbf{II}\|_{2}=a$. Let $\mathrm{vol}(M)$ and $\mathrm{vol}_{c}(M)$ denote the Euclidean volume of $M$ and the Riemannian volume of $(M,g_{c})$, respectively. Then
$\mathrm{vol}_{c}(M)\;\leq\;\Bigl(1+c\Bigl(\sup_{p\in M}\;\sup_{\begin{subarray}{c}v\in T_{p}M\\ \|v\|_{2}=1\end{subarray}}\|\mathbf{II}_{p}(v,\cdot)\|_{\operatorname{HS}}\Bigr)^{2}\Bigr)^{d/2}\mathrm{vol}(M).$
In particular,
$\mathrm{vol}_{c}(M)\leq(1+ca^{2}\min(d,D-d))^{d/2}\mathrm{vol}(M).$
Proof.
By Lemma 3.1 and the fact that $\mathbf{Gr}^{+}(D,d)\to\mathbf{Gr}(D,d)$ is a local isometry, for every $p\in M$ and $X\in\mathfrak{X}(M)$,
$(g_{c})_{p}(X,X)=\|X_{p}\|_{2}^{2}+c\,\|\mathbf{II}_{p}(X_{p},\cdot)\|_{\operatorname{HS}}^{2}.$
Set $K=\sup_{p\in M}\;\sup_{\begin{subarray}{c}v\in T_{p}M\\ \|v\|_{2}=1\end{subarray}}\|\mathbf{II}_{p}(v,\cdot)\|_{\operatorname{HS}}$. Then
$(g_{c})_{p}(X,X)\;\leq\;(1+cK^{2})\,\|X_{p}\|_{2}^{2}.$
Hence, the associated volume forms satisfy
$d\mathrm{vol}_{c}\;\leq\;(1+cK^{2})^{d/2}\,d\mathrm{vol}.$
Integrating over $M$ yields the first inequality. The second inequality follows from
$\|\mathbf{II}_{p}(X_{p},\cdot)\|_{\operatorname{HS}}^{2}\;\leq\;\min(d,D-d)\,\|\mathbf{II}_{p}\|_{2}^{2}\|X_{p}\|_{2}^{2}\;\leq\;\min(d,D-d)\,a^{2}\|X_{p}\|_{2}^{2}.$
3.2 Normal curvature reduction
Lemma 3.2.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable $d$-dimensional submanifold with normal bundle $\nu$. Denote by $\nu_{c}$ the normal bundle of $(M,g_{c})\subset\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$. Then for each $p\in M$,
$(\nu_{c})_{p}=\bigl\{(w,0)+(-cS_{p}(Q),Q)\;\big|\;w\in\nu_{p},\;Q\in T_{\bar{\mathbf{g}}_{c}^{+}(p)}\mathbf{Gr}_{c}^{+}(D,d)\bigr\}\subset T_{(p,\bar{\mathbf{g}}_{c}^{+}(p))}(\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)),$
where $S_{p}(Q)\in T_{p}M$ is the unique vector characterized by $\langle v,S_{p}(Q)\rangle_{\mathbb{R}^{D}}=\langle\mathbf{II}_{p}(v,\cdot),Q\rangle_{\operatorname{HS}}$ for every $v\in T_{p}M$.
Proof.
Fix $p\in M$. For every $(X,Q)\in(\nu_{c})_{p}$ and $v\in T_{p}M$, we have
$\langle v,X\rangle_{\mathbb{R}^{D}}+c\,\langle\mathbf{II}_{p}(v,\cdot),Q\rangle_{\operatorname{HS}}=0.$
Write $X=X^{\top}+X^{\perp}$ where $X^{\top}\in T_{p}M$ and $X^{\perp}\in\nu_{p}$. Since $\langle v,X^{\perp}\rangle_{\mathbb{R}^{D}}=0$, the condition reduces to $\langle v,X^{\top}\rangle_{\mathbb{R}^{D}}=-c\,\langle\mathbf{II}_{p}(v,\cdot),Q\rangle_{\operatorname{HS}}$ for every $v\in T_{p}M$, and letting $X^{\top}=-cS_{p}(Q)$ leads to the desired equality. The existence and uniqueness of the vector $S_{p}(Q)$ follow from the Riesz representation theorem. ∎
Lemma 3.3.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable $d$-dimensional submanifold, and let $X,Y\in\mathfrak{X}(M)$. Then
$\nabla^{\mathbf{Gr}_{c}^{+}(D,d)}_{d\bar{\mathbf{g}}_{c}^{+}(X)}\bigl(d\bar{\mathbf{g}}_{c}^{+}(Y)\bigr)=(\nabla^{M}_{X}\mathbf{II})(Y,\cdot)+\mathbf{II}(\nabla^{M}_{X}Y,\cdot)\in\operatorname{Hom}(TM,\nu),$
where $\nabla^{M}$ denotes the Levi–Civita connection on $M$.
Proof.
Since the covering $\mathbf{Gr}^{+}(D,d)\to\mathbf{Gr}(D,d)$ is a local isometry, we have
$\nabla^{\mathbf{Gr}_{c}^{+}(D,d)}_{d\bar{\mathbf{g}}_{c}^{+}(X)}\bigl(d\bar{\mathbf{g}}_{c}^{+}(Y)\bigr)=\nabla^{\mathbf{Gr}^{+}(D,d)}_{d\bar{\mathbf{g}}^{+}(X)}\bigl(d\bar{\mathbf{g}}^{+}(Y)\bigr)=\nabla^{\mathbf{Gr}(D,d)}_{\,d\bar{\mathbf{g}}(X)}\bigl(d\bar{\mathbf{g}}(Y)\bigr).$
By Lemma 3.1 and the Leibniz rule for tensor fields,
$\nabla^{\mathbf{Gr}(D,d)}_{d\bar{\mathbf{g}}(X)}\bigl(d\bar{\mathbf{g}}(Y)\bigr)=\nabla^{M}_{X}\bigl(\mathbf{II}(Y,\cdot)\bigr)=(\nabla^{M}_{X}\mathbf{II})(Y,\cdot)+\mathbf{II}(\nabla^{M}_{X}Y,\cdot),$
which proves the claim. ∎
Proof of Theorem 1.1.
Denote by $\nabla^{c}$ and $\nu_{c}$ the Levi–Civita connection on $\mathbb{R}^{D}\times\mathbf{Gr}^{+}_{c}(D,d)$ and the normal bundle of $(M,g_{c})\subset\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$, respectively. Suppose that there exists a unit local section (wcS(Q),Q)(w-cS(Q),Q) of $\nu_{c}$ near $p$ as in Lemma 3.2 such that
$\|w\|_{2}^{2}+c^{2}\|S(Q)\|_{2}^{2}+c\|Q\|_{\operatorname{HS}}^{2}=1,$
where $w$ is a section of $\nu$. By Lemma 3.3,
$\nabla^{c}_{(v,d\bar{\mathbf{g}}_{c}^{+}(v))}(v,d\bar{\mathbf{g}}_{c}^{+}(v))=\bigl(\nabla_{v}^{\mathbb{R}^{D}}v,\,\nabla^{\mathbf{Gr}_{c}^{+}(D,d)}_{d\bar{\mathbf{g}}_{c}^{+}(v)}d\bar{\mathbf{g}}_{c}^{+}(v)\bigr)=\bigl(\nabla_{v}^{M}v+\mathbf{II}(v,v),\,(\nabla_{v}^{M}\mathbf{II})(v,\cdot)+\mathbf{II}(\nabla_{v}^{M}v,\cdot)\bigr).$
Since
$d(\iota\times\bar{\mathbf{g}}_{c}^{+})(\nabla_{v}^{M}v)=\bigl(\nabla_{v}^{M}v,\,\mathbf{II}(\nabla_{v}^{M}v,\cdot)\bigr),$
the projections of $\nabla^{c}_{(v,d\bar{\mathbf{g}}_{c}^{+}(v))}(v,d\bar{\mathbf{g}}_{c}^{+}(v))$ and $\bigl(\mathbf{II}(v,v),(\nabla_{v}^{M}\mathbf{II})(v,\cdot)\bigr)$ onto $\nu_{c}$ coincide. On the other hand,
$g^{\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)}\bigl((w,0),(\mathbf{II}(v,v),(\nabla_{v}^{M}\mathbf{II})(v,\cdot))\bigr)=\bigl\langle w,\mathbf{II}(v,v)\bigr\rangle_{\mathbb{R}^{D}},$
and since $S(Q)\perp\mathbf{II}(v,v)$,
$g^{\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)}\bigl((-cS(Q),Q),(\mathbf{II}(v,v),(\nabla_{v}^{M}\mathbf{II})(v,\cdot))\bigr)=c\bigl\langle Q,(\nabla_{v}^{M}\mathbf{II})(v,\cdot)\bigr\rangle_{\operatorname{HS}}.$
Therefore,
$|g^{\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)}((w-cS(Q),Q),(\mathbf{II}({v},{v}),(\nabla_{v}^{M}\mathbf{II})({v},\cdot)))|$ $=|\langle w,\mathbf{II}(v,v)\rangle_{\mathbb{R}^{D}}+c\langle Q,(\nabla_{v}^{M}\mathbf{II})({v},\cdot)\rangle_{\operatorname{HS}}|$
$\leq(\kappa(v)\|w\|_{2}+c\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}\|Q\|_{\operatorname{HS}})$
$\leq\sqrt{\kappa(v)^{2}+c\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}}\sqrt{\|w\|_{2}^{2}+c\|Q\|_{\operatorname{HS}}^{2}}$
$\leq\sqrt{\kappa(v)^{2}+c\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}}.$
where we used $\|w\|_{2}^{2}+c^{2}\|S(Q)\|_{2}^{2}+c\|Q\|_{\operatorname{HS}}^{2}=1$ in the last inequality. Hence
$\kappa_{c}(v)\leq\frac{\sqrt{\kappa(v)^{2}+c\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}}}{\|(v,d\bar{\mathbf{g}}_{c}^{+}(v))\|_{2}^{2}}=\frac{\sqrt{\kappa(v)^{2}+c\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}}}{1+c\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}.$
By the volume assumption in Theorem 1.1, we obtain
$\kappa_{c}(v)\,(\mathrm{vol}_{c}(M))^{1/d}$ $\leq\frac{\sqrt{\kappa(v)^{2}+c\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}}}{\sqrt{1+c\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}}\,(\mathrm{vol}(M))^{1/d}.$
We compute
$\frac{d}{dc}\left(\frac{\kappa(v)^{2}+c\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}}{1+c\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}\right)$ $=\frac{\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}-\kappa(v)^{2}\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}{(1+c\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2})^{2}}<0,$
by the assumption. This completes the proof. ∎
Remark 3.1.
In codimension one, if $\{e_{i}\}_{i=1}^{D-1}$ is an orthonormal eigenbasis of $T_{q}M$ with principal curvatures $\lambda_{1}(q),\dots,\lambda_{D-1}(q)$, then in this basis the metric $g_{c}$ satisfies
$(g_{c})_{q}]=I+cH_{q},\qquad H_{q}=\operatorname{diag}(\lambda_{1}(q)^{2},\dots,\lambda_{D-1}(q)^{2}),$
so that $\operatorname{tr}H_{q}=\sum_{i=1}^{D-1}\lambda_{i}(q)^{2}$. Therefore, for small $c>0$, we obtain
$d\mathrm{vol}_{c}=\sqrt{\det(I+cH_{q})}\,d\mathrm{vol}=\Bigl(1+\frac{c}{2}\,\sum_{i=1}^{D-1}\lambda_{i}(q)^{2}+O(c^{2})\Bigr)\,d\mathrm{vol}.$
In particular, regions where many principal curvatures are simultaneously large contribute most to the first-order increase of $\mathrm{vol}_{c}(M)$. On the other hand, for any $p\in M$ and any unit vector $v\in T_{p}M$, the volume condition in Theorem 1.1 states
$\mathrm{vol}_{c}(M)\;\leq\;\bigl(1+c\,\|\mathbf{II}_{p}(v,\cdot)\|_{\operatorname{HS}}^{2}\bigr)^{(D-1)/2}\mathrm{vol}(M)=\Bigl(1+\frac{c\,(D-1)}{2}\|\mathbf{II}_{p}(v,\cdot)\|_{\operatorname{HS}}^{2}+O(c^{2})\Bigr)\,\mathrm{vol}(M),$
Thus, for small $c$, one may interpret this condition as requiring that at points $p$ and directions $v$ the normal curvature is significantly larger than the average curvature of $M$. In higher codimension, the analogous condition may require such a largeness assumption for all sections of the normal bundle.
Remark 3.2.
The one-parameter family of metrics $\{g_{c}\}_{c>0}$ may be viewed as an extrinsic curvature-regularizing deformation of $M$ at scale $c$. Recall that $\mathbf{II}_{c}$ denotes the second fundamental form of $(M,g_{c})\subset\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$. For $p\in M$ and a unit vector $v\in T_{p}M$, the normal curvature $\kappa_{c}(v)$ relative to $\mathrm{vol}_{c}(M)$ decreases as $c$ grows by Theorem 1.1 if $\kappa(v)$ is sufficiently large. Thus, the deformation $g_{c}$ can reduce $\|\mathbf{II}_{c}\|_{2}$ relative to $\mathrm{vol}_{c}(M)$ for appropriate $c>0$.
The condition $\kappa(v)>\frac{\|\nabla_{v}^{M}(\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}}{\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}}$ indicates that the normal curvature of $M$ in the direction $v$ dominates the logarithmic variation of the second fundamental form.
Theorem 3.2.
Let $M\subset\mathbb{R}^{D}$ be a smooth $d$-dimensional submanifold. For a unit-speed geodesic $\gamma\subset M$ such that $\gamma^{\prime}(t)=v(t)$, suppose that the tensor $\mathbf{II}_{\gamma(t)}(v(t),\cdot)$ does not vanish along $\gamma$. Then
$\left|\frac{d}{dt}\log\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}\right|\leq\frac{\bigl\|\nabla^{M}_{v}(\mathbf{II})(v,\cdot)\bigr\|_{\operatorname{HS}}}{\bigl\|\mathbf{II}(v,\cdot)\bigr\|_{\operatorname{HS}}}.$
The proof is given in Appendix A.
3.3 Extension of the width of bottlenecks
Recall that for an embedded submanifold $M\subset\mathbb{R}^{D}$, we define
$L(M)=\min_{(q_{1},q_{2})\ \text{bottleneck}}\frac{1}{2}\,\|q_{1}-q_{2}\|_{\mathbb{R}^{D}},\qquad L_{c}(M)=\min_{(q_{1},q_{2})\ \text{bottleneck}}\frac{1}{2}\,d_{c}(q_{1},q_{2}).$
By definition, we have $L_{c}(M)\geq L(M)$. In general there is no meaningful lower bound for $L_{c}(M)$ that improves on $L(M)$ without additional assumptions on $M$. However, for closed hypersurfaces with small normal curvatures, we obtain nontrivial lower bounds for $L_{c}(M)$ that are strictly larger than $L(M)$.
Lemma 3.4.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable hypersurface with a unit normal vector field $\nu$. Let $q_{1},q_{2}\in M$ be two distinct points such that the open line segment between $q_{1}$ and $q_{2}$ is disjoint from $M$, and suppose that the vectors $\nu_{q_{1}}$ and $\nu_{q_{2}}$ are both parallel to the line through $q_{1}$ and $q_{2}$. Then the Gauss map $\bar{\mathbf{g}}^{+}\colon M\to S^{D-1}$ sends $q_{1}$ and $q_{2}$ to antipodal points on $S^{D-1}$.
Proof.
Since the open line segment between $q_{1}$ and $q_{2}$ does not intersect $M$, the Jordan–Brouwer separation theorem implies that it is contained entirely either in $\mathbb{R}^{D}\setminus\mathrm{int}(M)$ or in $\mathrm{int}(M)$. If the segment lies inside $\mathrm{int}(M)$, then the outward normal at $q_{1}$ points in the direction of $\overrightarrow{q_{1}q_{2}}$ while the outward normal at $q_{2}$ points in the opposite direction $\overrightarrow{q_{2}q_{1}}$. If instead the segment lies outside $M$, the same conclusion holds. Hence $\nu_{q_{1}}=-\nu_{q_{2}}$, and the Gauss map $\bar{\mathbf{g}}^{+}$ sends $q_{1}$ and $q_{2}$ to antipodal points on $S^{D-1}$. ∎
Theorem 3.3.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable hypersurface. Suppose that the reach of $M\subset\mathbb{R}^{D}$ is attained at a bottleneck of $M$ (see the second case in Theorem 2.2) with $\textsf{rch}_{\mathbb{R}^{D}}(M)=L(M)$. Then
$L_{c}(M)\geq\min\!\left(\tfrac{1}{2}\sqrt{\,4L(M)^{2}+c\pi^{2}\,},\,2L(M)\right).$
Proof.
Let $\nu$ denote the unit normal vector field of $M\subset\mathbb{R}^{D}$ and suppose that $q_{1},q_{2}\in M$ form a bottleneck of $M\subset\mathbb{R}^{D}$. If $\nu_{q_{1}}=\nu_{q_{2}}$ (see Figure 3), then by Lemma 3.4, the line segment $\overline{q_{1}q_{2}}$ contains another point $q_{3}\in M$. We will show that $\|q_{1}-q_{3}\|_{\mathbb{R}^{D}}\geq 2L(M)$.
Suppose for contradiction that $\|q_{1}-q_{3}\|_{\mathbb{R}^{D}}<2L(M)$ and let the midpoint $q_{4}=\frac{q_{1}+q_{3}}{2}$. Then
$\|q_{1}-q_{4}\|_{\mathbb{R}^{D}}<L(M),\quad\|q_{3}-q_{4}\|_{\mathbb{R}^{D}}<L(M),\quad\overline{q_{1}q_{4}}\perp T_{q_{1}}M.$
Using the distance function $\|q_{4}-\cdot\|_{\mathbb{R}^{D}}$ with the initial point $q_{3}$, one can find $q_{5}\in M$ such that $\|q_{5}-q_{4}\|_{\mathbb{R}^{D}}<L(M)$ and $T_{q_{5}}M\perp\overline{q_{5}q_{4}}.$ If $q_{5}=q_{1}$, then there exists a curve $\delta$ joining $q_{3}$ and $q_{1}$ such that $\|\delta(t)-q_{4}\|_{\mathbb{R}^{D}}$ is decreasing. However, since $\|q_{3}-q_{4}\|_{\mathbb{R}^{D}}=\|q_{4}-q_{1}\|_{\mathbb{R}^{D}}$, the function $\|\delta(t)-q_{4}\|_{\mathbb{R}^{D}}$ must be constant, yielding a contradiction. So $q_{1}\neq q_{5}$. Then the point $q_{4}$ satisfies
$\|q_{1}-q_{4}\|_{\mathbb{R}^{D}}<L(M),\quad\|q_{5}-q_{4}\|_{\mathbb{R}^{D}}<L(M),\quad T_{q_{1}}M\perp\overline{q_{1}q_{4}},\quad T_{q_{5}}M\perp\overline{q_{5}q_{4}}.$
Since the reach of $M\subset\mathbb{R}^{D}$ equals $L(M)$, any point $y$ with $d_{\mathbb{R}^{D}}(y,M)<L(M)$ has a unique nearest point $p\in M$ such that $y-p\perp T_{p}M$ and $\|y-p\|_{\mathbb{R}^{D}}=d_{\mathbb{R}^{D}}(y,M)$. For $y=q_{4}$ we obtain two such points $q_{1}$ and $q_{5}$, which is a contradiction. Therefore, we conclude $\|q_{1}-q_{3}\|_{\mathbb{R}^{D}}\geq 2L(M)$ and one can deduce $\|q_{2}-q_{3}\|_{\mathbb{R}^{D}}\geq 2L(M)$ in the same way. Hence $\|q_{1}-q_{2}\|_{\mathbb{R}^{D}}=\|q_{1}-q_{3}\|_{\mathbb{R}^{D}}+\|q_{3}-q_{2}\|_{\mathbb{R}^{D}}\geq 4L(M)$.
$q_{1}$$q_{3}$$q_{2}$$q_{5}$$q_{4}$
Figure 3: Geometric configuration in the proof of Theorem 3.3.
If $\nu_{q_{1}}=-\nu_{q_{2}}$, then $\|q_{1}-q_{2}\|_{\mathbb{R}^{D}}\geq 2L(M)$, and by Lemma 3.4, the geodesic distance $d_{g^{\mathbf{Gr}^{+}}}(\bar{\mathbf{g}}(q_{1}),\bar{\mathbf{g}}(q_{2}))$ equals the length of the half of the great circle in $S^{D-1}$. Hence $d_{c}(q_{1},q_{2})\geq\sqrt{(2L(M))^{2}+c\cdot(\pi)^{2}}$. Considering both cases together, we conclude that
$L_{c}(M)\geq\min\!\left(\tfrac{1}{2}\sqrt{\,4L(M)^{2}+c\pi^{2}\,},\,2L(M)\right).$
Proof of Theorem 1.2.
Let $a=\|\mathbf{II}\|_{2}>0$. Since $L(M)\leq\frac{1}{a}$, Theorem 2.2 implies that the reach of $M\subset\mathbb{R}^{D}$ is attained at a bottleneck. By Theorem 3.3, for $c\in[0,\frac{12L(M)^{2}}{\pi^{2}}]$,
$L_{c}(M)\geq\frac{1}{2}\sqrt{4L(M)^{2}+c\pi^{2}}.$
On the other hand, by Theorem 3.1,
$\mathrm{vol}_{c}(M)\leq(1+ca^{2})^{(D-1)/2}\mathrm{vol}(M).$
Combining these inequalities, we obtain
$\frac{L_{c}(M)}{\mathrm{vol}_{c}(M)^{1/(D-1)}}\geq\frac{1}{2(\mathrm{vol}(M))^{1/(D-1)}}\sqrt{\frac{4L(M)^{2}+c\pi^{2}}{1+ca^{2}}}.$
Since $L(M)\leq\frac{1}{a}$ again, we compute
$\frac{d}{dc}\left(\frac{4L(M)^{2}+c\pi^{2}}{1+ca^{2}}\right)=\frac{\pi^{2}-4L(M)^{2}a^{2}}{(1+ca^{2})^{2}}>0,$
which completes the proof. ∎
As an example, we consider the standard embedded torus $T^{2}\subset\mathbb{R}^{3}$, for which the systole and the minima width of bottlenecks are small compared to its volume. We show that under the embedding into $\mathbb{R}^{3}\times\mathbf{Gr}_{c}^{+}(3,2)$, the normalized quantities $\mathrm{sys}_{c}(T^{2})^{2}/\mathrm{vol}_{c}(T^{2})$ and $L_{c}(T^{2})^{2}/\mathrm{vol}_{c}(T^{2})$ can be significantly increased. For comparison, recall Loewner’s torus inequality [pu1952some] in the classical setting, stating that every Riemannian $2$-torus $\mathbb{T}^{2}$ equipped with an arbitrary metric satisfies
$\frac{\mathrm{sys}(\mathbb{T}^{2})^{2}}{\mathrm{vol}(\mathbb{T}^{2})}\leq\frac{2}{\sqrt{3}}.$
Example 1.
Consider the standard embedded torus $T^{2}\subset\mathbb{R}^{3}$ parametrized by
$(x,y,z)=\bigl((R+r\cos v)\cos u,\ (R+r\cos v)\sin u,\ r\sin v\bigr),\qquad u,v\in[0,2\pi),\ 0<r<R.$
Assume that $\frac{r}{R}\leq\frac{1}{m}$ for some $m\geq 2$, and set $c=(R-r)^{2}$. For every $\varepsilon>0$ there exists $M=M(\varepsilon)>0$ such that, if $m>M$, then
$\frac{\mathrm{sys}_{c}(T^{2})^{2}}{\mathrm{vol}_{c}(T^{2})}>\frac{2\pi}{2\pi+4}-\varepsilon,\qquad\frac{L_{c}(T^{2})^{2}}{\mathrm{vol}_{c}(T^{2})}>\frac{1}{2\pi(2\pi+4)}-\varepsilon,$
where $\mathrm{sys}_{c}(T^{2})$ and $\mathrm{vol}_{c}(T^{2})$ denote the systole and the Riemannian volume of $(T^{2},g_{c})$, respectively. In particular, as $\frac{r}{R}\to 0$, the normalized quantities $\mathrm{sys}_{c}(T^{2})^{2}/\mathrm{vol}_{c}(T^{2})$ and $L_{c}(T^{2})^{2}/\mathrm{vol}_{c}(T^{2})$ approach the positive constants $\frac{2\pi}{2\pi+4}$ and $\frac{1}{2\pi(2\pi+4)}$, respectively. The proof is deferred to Appendix B.
3.4 Range of radii for homotopy equivalences
Lemma 3.5 ([albano2015cut, Theorem 3.2], [warner1966extension, Corollary 4.2]; modified).
Let $S$ be a smooth submanifold of a closed Riemannian manifold $N$. Assume that the sectional curvature of $N$ is everywhere bounded above by $K>0$ and that the operator norm of the second fundamental form of $S\subset N$ is bounded above by $\Lambda>0$. Let $t_{\mathrm{foc}}$ denote the first focal time of $S$ (see the first case in Figure 2). Then
$t_{\mathrm{foc}}\;\geq\;\frac{1}{\sqrt{K}}\arctan\!\left(\frac{\sqrt{K}}{\Lambda}\right).$
Theorem 3.4.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable $d$-dimensional submanifold. Suppose that for a point $q\in M$, there exist distinct unit-speed geodesics $(\gamma_{1},\Gamma_{1})$ and $(\gamma_{2},\Gamma_{2})$ in $\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$ satisfying
$(\gamma_{1},\Gamma_{1})(0)=(\gamma_{2},\Gamma_{2})(0)=(q,\bar{\mathbf{g}}^{+}_{c}(q)),\qquad(\gamma_{1},\Gamma_{1})(T)=(\gamma_{2},\Gamma_{2})(T).$
Then
$T\;\geq\;\frac{\sqrt{c}\,\pi}{2}.$
Proof.
The Euclidean space $\mathbb{R}^{D}$ with the standard metric has infinite conjugate radius. By [bendokat2024grassmann, Theorem 7.2], the Grassmannian $\mathbf{Gr}(D,d)$ has conjugate radius greater than $\frac{\pi}{2}$. Since the oriented Grassmannian $\mathbf{Gr}^{+}(D,d)$ is a connected Riemannian covering of $\mathbf{Gr}(D,d)$, its conjugate radius is also greater than $\frac{\pi}{2}$. Under the scaled metric on $\mathbf{Gr}^{+}_{c}(D,d)$, the conjugate radius is therefore greater than $\frac{\sqrt{c}\,\pi}{2}$. Hence, along any unit-speed geodesic in $\mathbb{R}^{D}\times\mathbf{Gr}^{+}_{c}(D,d)$ there is no conjugate point before time $\frac{\sqrt{c}\,\pi}{2}$. Thus, the existence of two distinct unit-speed geodesics with common endpoints at time $T$ implies $T\;\geq\;\frac{\sqrt{c}\,\pi}{2}$. ∎
Proof of Theorem 1.3.
Let $\mathsf{rch}_{\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)}(M,g_{c})=T$. Since the sectional curvature of the Grassmannian $(\mathbf{Gr}(D,d),g^{\mathbf{Gr}(D,d)})$ is bounded above by $2$ [bendokat2024grassmann, Proposition 4.1], the sectional curvature of the product manifold $\mathbb{R}^{D}\times\mathbf{Gr}^{+}_{c}(D,d)$ is bounded above by $\tfrac{2}{c}$. By Theorem 3.5, if $T>0$ is a first focal time at some point $p\in M$, then
$T\;\geq\;\sqrt{\frac{c}{2}}\,\arctan\sqrt{\frac{2}{c\|\mathbf{II}_{c}\|_{2}^{2}}}.$
If there exist two distinct unit-speed distance-minimal geodesics between the same pair of points with length $T$, then by Theorem 3.4, we have $T\geq\frac{\sqrt{c}\pi}{2}$. Otherwise, there exist points $q_{1},q_{2}\in M$ and a point $(p,P)\in\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$, together with unit-speed distance-minimal geodesics $(\gamma_{1},\Gamma_{1})$ and $(\gamma_{2},\Gamma_{2})$ from $(q_{1},\bar{\mathbf{g}}_{c}^{+}(q_{1}))$ and $(q_{2},\bar{\mathbf{g}}_{c}^{+}(q_{2}))$ to $(p,P)$, respectively, each of length $T$. Since $\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$ is complete, by [singh1987cut, Theorem 1], the two geodesics meet at $(p,P)$ at angle $\pi$ and $(p,P)$ is one of the midpoints of the geodesic segment joining $(q_{1},\bar{\mathbf{g}}_{c}^{+}(q_{1}))$ and $(q_{2},\bar{\mathbf{g}}_{c}^{+}(q_{2}))$. By the first variational principle, the two geodesics $(\gamma_{1},\Gamma_{1})$ and $(\gamma_{2},\Gamma_{2})$ are orthogonal to $T_{(q_{1},\bar{\mathbf{g}}_{c}^{+}(q_{1}))}(M,g_{c})$ and $T_{(q_{2},\bar{\mathbf{g}}_{c}^{+}(q_{2}))}(M,g_{c})$, respectively. Therefore $T\geq L_{c}^{\prime}(M)$ by the definition of $L_{c}^{\prime}(M)$.
Combining these cases and using Theorem 2.3, we obtain
$T\;\geq\;\min\!\left(\sqrt{\frac{c}{2}}\,\arctan\sqrt{\frac{2}{c\|\mathbf{II}_{c}\|_{2}^{2}}},\;\frac{\sqrt{c}\,\pi}{2},\;L_{c}^{\prime}(M)\right).$
By the nerve lemma and Theorem 2.4, it follows that for every
$0<r<\min\!\left(\sqrt{\frac{c}{2}}\,\arctan\sqrt{\frac{2}{c\|\mathbf{II}_{c}\|_{2}^{2}}},\;\frac{\sqrt{c}\,\pi}{2},\;{L_{c}}^{\prime}(M)\right),$
there are homotopy equivalences
$\check{C}(M,\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d);r)\;\simeq\;\bigcup_{q\in M}B_{\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)}\bigl((q,\bar{\mathbf{g}}_{c}^{+}(q)),r\bigr)\;\simeq\;M,$
which completes the proof. ∎
Remark 3.3.
Theorem 1.1 together with Remark 3.2 suggests that, for suitable choices of $c>0$, the directions with large normal curvature are damped so that $\|\mathbf{II}_{c}\|_{2}$ can decrease relative to $\mathrm{vol}_{c}(M)$. In view of Theorem 1.3, this suggests that for such $c$ the first term in the lower bound for $T$ can become larger relative to the volume, compared to the Euclidean metric.
Remark 3.4.
Theorem 1.2 shows that, in the case for which the reach of an orientable hypersurface $M\subset\mathbb{R}^{D}$ is realized by a bottleneck, the distance between such bottleneck pairs normalized by the volume grows. According to Lemma 3.2, there is no direct relation between $L_{c}(M)$ and $L_{c}^{\prime}(M)$ in general, since the normal directions at $p\in M$ may tilt when $M$ is embedded into $\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$. However, Lemma 3.2 also implies that this tilting is controlled by $c\|\mathbf{II}\|_{\operatorname{HS}}$. Thus, when $c\|\mathbf{II}\|_{\operatorname{HS}}$ is small, the normals are only slightly tilted and it is reasonable to expect $L_{c}(M)$ and $L_{c}^{\prime}(M)$ to be close to each other, so that the bound for the width of bottlenecks of the hypersurface $M$ from Theorem 1.2 remains valid as resolving such neck structures under this embedding.
Remark 3.5.
We provide a range of radii $r$ where the ambient Čech complex is homotopy equivalent to the underlying manifold $M$, according to Theorem 2.4. By Theorem 2.5, the same conclusion holds for the ambient Vietoris–Rips complexes $\mathrm{VR}(\mathbf{Y},\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d);r)$ for any sufficiently dense subset $\mathbf{Y}\subset M$.
3.5 Stability
To prove Theorem 1.4, we use the following perturbation bound for orthogonal projectors.
Theorem 3.5 ([stewart1998perturbation, Theorem 4]; modified).
[…]
$\|P_{\widetilde{A}}-P_{A}\|_{\operatorname{HS}}\leq C\,\frac{\|E\|_{\operatorname{HS}}}{\sigma_{\min}(A)}.$
Proof of Theorem 1.4.
Let $F_{i}:M\to\mathbb{R}^{D}$ be smooth embeddings with $M_{i}=F_{i}(M)\subset\mathbb{R}^{D}$ for each $i\in\mathbb{N}$. Assume that there exists an embedded manifold $M_{\infty}=F_{\infty}(M)\subset\mathbb{R}^{D}$ such that $F_{i}\rightarrow F_{\infty}$ in the Whitney $C^{1}$ topology. Since $C^{1}$ convergence implies that $M_{i}$ converges to $M_{\infty}$ in the Hausdorff distance and $M_{\infty}$ is compact, there exists a compact subset $K\subset\mathbb{R}^{D}$ containing both $M_{\infty}$ and $M_{i}$ for every $i\in\mathbb{N}$.
Choose finitely many coordinate charts $\{(U_{\alpha},\varphi_{\alpha})\}_{\alpha=1}^{m}$ on $M$ and coordinate charts $\{(V_{\alpha},\psi_{\alpha})\}_{\alpha=1}^{m}$ on $K$ such that $F_{\infty}(U_{\alpha})\subset V_{\alpha}$ for each $\alpha$. By the Whitney $C^{1}$ convergence $F_{i}\to F_{\infty}$, for each $\alpha$ and each $\varepsilon_{\alpha}>0$ there exists $n_{\alpha}\in\mathbb{N}$ such that for every $i\geq n_{\alpha}$,
$\sup_{x\in\varphi_{\alpha}(U_{\alpha})}\big\|(\psi_{\alpha}\circ F_{i}\circ\varphi_{\alpha}^{-1})(x)-(\psi_{\alpha}\circ F_{\infty}\circ\varphi_{\alpha}^{-1})(x)\big\|_{\mathbb{R}^{D}}<\varepsilon_{\alpha},$
and
$\sup_{x\in\varphi_{\alpha}(U_{\alpha})}\big\|d(\psi_{\alpha}\circ F_{i}\circ\varphi_{\alpha}^{-1})(x)-d(\psi_{\alpha}\circ F_{\infty}\circ\varphi_{\alpha}^{-1})(x)\big\|_{\operatorname{HS}}<\varepsilon_{\alpha}.$
Let $n=\max\{n_{\alpha}\}_{1\leq\alpha\leq m}$ and $\varepsilon=\varepsilon(n)=\min\{\varepsilon_{\alpha}\}_{1\leq\alpha\leq m}$. Since $F_{\infty}:M\to K$ is an embedding and $M_{\infty}$ is compact, the differential $d(\psi_{\alpha}\circ F_{\infty}\circ\varphi_{\alpha}^{-1})(x)$ has full rank and depends continuously on $x$. Hence there exists a constant $\delta>0$ such that, for every $\alpha$ and every $x\in\varphi_{\alpha}(U_{\alpha})$, the smallest singular value of $d(\psi_{\alpha}\circ F_{\infty}\circ\varphi_{\alpha}^{-1})(x)$ is at least $\delta$.
Denote the principal angles between $\bar{\mathbf{g}}(F_{i}(q))$ and $\bar{\mathbf{g}}(F_{\infty}(q))$ by $\theta_{1}(i,q)\geq\dots\geq\theta_{d}(i,q)$ for $i\geq n$ and $q\in M$. For a dd–dimensional subspace $W\subset\mathbb{R}^{D}$ we write $P_{W}$ for the orthogonal projector onto $W$. By Theorem 3.5 and the identity $\bigl\|P_{\bar{\mathbf{g}}(F_{i}(q))}-P_{\bar{\mathbf{g}}(F_{\infty}(q))}\bigr\|_{\operatorname{HS}}^{2}=2\sum_{j=1}^{d}\sin^{2}\theta_{j}(i,q)$, we obtain
 $\sqrt{\,2\sum_{j=1}^{d}\sin^{2}\theta_{j}(i,q)\,}<\frac{C}{\delta}\varepsilon$
for some constant $C>0$. Hence
$d_{\mathbf{Gr}(D,d)}\!\bigl(\bar{\mathbf{g}}(F_{i}(q)),\bar{\mathbf{g}}(F_{\infty}(q))\bigr)=\sqrt{\sum_{j=1}^{d}\theta_{j}^{2}(i,q)}\;\leq\;\frac{\pi}{2}\sqrt{\sum_{j=1}^{d}\sin^{2}\theta_{j}(i,q)}\;<\;\frac{C\pi}{2\sqrt{2}\delta}\varepsilon.$
Take $n$ large enough so that $\varepsilon<\tfrac{\sqrt{2}\delta}{C}$. Suppose that there exist $x_{1},x_{2}\in\varphi_{\alpha}(U_{\alpha})$ such that
$\det(d\big(\psi_{\alpha}\!\circ\!F_{i}\!\circ\!\varphi_{\alpha}^{-1}\big)(x_{1})^{\top}d\big(\psi_{\alpha}\!\circ\!F_{i}\!\circ\!\varphi_{\alpha}^{-1}\big)(x_{1}))>0$
but
$\det(d\big(\psi_{\alpha}\!\circ\!F_{i}\!\circ\!\varphi_{\alpha}^{-1}\big)(x_{2})^{\top}d\big(\psi_{\alpha}\!\circ\!F_{i}\!\circ\!\varphi_{\alpha}^{-1}\big)(x_{2}))<0.$
Since the map
$x\longmapsto\det\!\Bigl(d(\psi_{\alpha}\circ F_{i}\circ\varphi_{\alpha}^{-1})(x)^{\top}d(\psi_{\alpha}\circ F_{i}\circ\varphi_{\alpha}^{-1})(x)\Bigr)$
is continuous on $\varphi_{\alpha}(U_{\alpha})$, there exists a point $x_{3}\in\varphi_{\alpha}(U_{\alpha})$ such that
$\det(d\big(\psi_{\alpha}\!\circ\!F_{i}\!\circ\!\varphi_{\alpha}^{-1}\big)(x_{3})^{\top}d\big(\psi_{\alpha}\!\circ\!F_{i}\!\circ\!\varphi_{\alpha}^{-1}\big)(x_{3}))=0,$
so that
$d_{\mathbf{Gr}(D,d)}(\bar{\mathbf{g}}(F_{i}(\varphi_{\alpha}(x_{3}))),\bar{\mathbf{g}}(F_{\infty}(\varphi_{\alpha}(x_{3}))))\geq\frac{\pi}{2},$
which contradicts the assumption that $\varepsilon<\tfrac{\sqrt{2}\delta}{C}$. Therefore, we may choose orientations for each $F_{i}(M)$ with
$\det(d\big(\psi_{\alpha}\!\circ\!F_{i}\!\circ\!\varphi_{\alpha}^{-1}\big)(x)^{\top}d\big(\psi_{\alpha}\!\circ\!F_{i}\!\circ\!\varphi_{\alpha}^{-1}\big)(x))>0$
for every $x\in\varphi_{\alpha}(U_{\alpha})$ and $i\geq n$, and hence
$d_{\mathbf{Gr}^{+}(D,d)}(\bar{\mathbf{g}}^{+}(F_{i}(q)),\bar{\mathbf{g}}^{+}(F_{\infty}(q)))=d_{\mathbf{Gr}(D,d)}(\bar{\mathbf{g}}(F_{i}(q)),\bar{\mathbf{g}}(F_{\infty}(q)))$
for every $q\in M$.
Therefore, for every $q\in M$ and $i\geq n$,
$d_{c}(F_{i}(q),F_{\infty}(q))<\varepsilon\sqrt{1+c\frac{C^{2}\pi^{2}}{8\delta^{2}}}.$
In particular,
$d^{\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)}_{H}(F_{i}(M),F_{\infty}(M))<\varepsilon\sqrt{1+c\frac{C^{2}\pi^{2}}{8\delta^{2}}}.$
Applying Theorem 2.1 completes the proof. ∎
4 Computational examples
The estimation of tangent spaces of an underlying smooth manifold via local PCA [kambhatla1997dimension, abdi2010principal] has recently emerged in TDA. In one approach, local PCA is used to estimate tangent-normal coordinates at each point, and the resulting ellipsoids are incorporated into a modified filtration metric [kalisnik2020finding, kalivsnik2024persistent]. On the other hand, local PCA has also been employed to approximate characteristic classes of vector bundles over the underlying manifold [scoccola2023approximate, bohlsen2025counting, turow2025discrete]. In [tinarrage2023recovering], locally estimated tangent spaces are used to recover the homology of immersed smooth manifolds with self-intersections. In [bauer2023cycling], a unit vector field of the flow was implemented for computing a filtration metric to detect a cycling behavior in a cubical setting.
We outline a procedure to estimate the distance $d_{c}$ in Definition 1.1 using local PCA. First, we choose a finite subset $\mathbf{X}\subset M$ of the smooth closed orientable $d$-dimensional manifold $M\subset\mathbb{R}^{D}$ to estimate its tangent bundle. Next, we select a subset $\mathbf{Y}\subset\mathbf{X}$ to build a Vietoris–Rips filtration. With neighborhood size chosen as $k\sim|\mathbf{X}|^{2/(n+2)}$, we perform local PCA at each point $x\in\mathbf{X}$ using its k-nearest neighbors to obtain an orthonormal basis for the estimated tangent space at $x$.
Theorem 4.1 ([singer2012vector, Theorem B.1]).
Let $\varepsilon=\varepsilon(n)$ be a sequence of positive numbers with $\varepsilon=O\bigl(n^{-\frac{1}{d+2}}\bigr)$. Assume that for each $x_{i}\in\mathbf{X}$, the set $\bigl\{y\in\mathbf{X}\mid d_{\mathbb{R}^{D}}(x_{i},y)\leq\varepsilon\bigr\}$ contains at least $d$ points, and let $\widehat{T_{x_{i}}M}$ be the tangent space at $x_{i}$ estimated from this neighborhood by local PCA. Then, with high probability,
$d_{\mathbf{Gr}(D,d)}\bigl(T_{x_{i}}M,\widehat{T_{x_{i}}M}\bigr)=O\bigl(\varepsilon^{\frac{3}{2}}\bigr)=O\bigl(n^{-\frac{3}{2(d+2)}}\bigr).$
For each $x\in M$, fix an ordered orthonormal basis for $T_{x}M$ compatible with the given orientation on $M$, and let $B_{x}$ be the matrix whose columns are this basis. Let $\widehat{T_{x}M}$ be the estimated tangent space at $x$, and choose an ordered orthonormal basis for $\widehat{T_{x}M}$ so that it induces the same orientation as $T_{x}M$. Denote the corresponding basis matrix by $\widehat{B_{x}}$. To establish that such orientations can be chosen consistently across tangent spaces, we prove the following theorem.
Theorem 4.2.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed connected orientable $d$-dimensional submanifold with $\mathsf{rch}_{\mathbb{R}^{D}}(M)=\tau>0$. If two points $p,q\in M$ satisfy $\|p-q\|_{\mathbb{R}^{D}}<\frac{\tau}{2}$, then
$\det(B_{p}^{\top}B_{q})>0.$
We postpone the proof to Appendix C. If two points $x_{1}$ and $x_{2}$ in $\mathbf{X}$ are adjacent and the basis matrix $\widehat{B_{x_{1}}}$ is oriented, then we orient $\widehat{B_{x_{2}}}$ by flipping the sign of its last row whenever $\det(\widehat{B_{x_{1}}}^{\top}\widehat{B_{x_{2}}})<0$. This yields a consistent orientation of the estimated tangent spaces over $\mathbf{X}$. We choose $c>0$ so that the diameters of $M$ and $\mathbf{Gr}_{c}^{+}(D,d)$ coincide. Approximating $\mathrm{diam}(M)$ by $\widehat{\mathrm{diam}}(M)=\mathrm{diam}(\mathbf{Y})$ and using $\mathrm{diam}(\mathbf{Gr}^{+}(D,d))=\max\bigl(\pi,\frac{\pi}{2}\,\sqrt{\min(d,D-d)}\bigr)$ ([kozlov2000geometry, Theorem 12.6]), we set
$c=\frac{\mathrm{diam}(\mathbf{Y})^{2}}{\max(\pi,\frac{\pi}{2}\sqrt{\min(d,D-d)})^{2}}.$
We then compute the distance matrix $\mathbf{D}$ with entries
$\bigl(\mathbf{D}\bigr)_{ij}=\sqrt{\|y_{i}-y_{j}\|_{\mathbb{R}^{D}}^{2}+c\,d_{\mathbf{Gr}^{+}(D,d)}(\widehat{T_{y_{i}}M},\widehat{T_{y_{j}}M})^{2}}$
for $1\leq i,j\leq|\mathbf{Y}|$ and perform Vietoris–Rips persistent homology with respect to $\mathbf{D}$. We compute persistent homology using the Ripser library [bauer2021ripser], with coefficients in $\mathbb{Z}/2$.
4.1 Time-delay embedding of the double-gyre attractor
The driven double-gyre system on $\psi$ in $(x,y)\in[0,2]\times[0,1]$ is specified by the stream function
$\phi(x,y,t)=C\sin(\pi f(x,t))\sin(\pi y),\quad f(x,t)=(\eta\sin(\omega t))x^{2}+(1-2\eta\sin(\omega t))x.$ (1)
This system was first introduced in [shadden2005definition] to describe geophysical flows in oceanic systems [coulliette2001intergyre, poje1999geometry]. For a continuous dynamical system $(M,\Phi)$, a compact invariant subset $A\subset M$ is called an attractor if there exists an open neighborhood $U\subset M$ of $A$ such that
$A=\bigcap_{t\geq 0}\{\Phi(p,t)\,:\,p\in U\}.$
For more details, see [smale1967differentiable] or [perea2019topological]. We choose $20000$ points from the system (1) with times uniformly distributed over $[0,10000]$. The parameters are $C=0.1$, $\eta=0.1$, $\omega=\pi/5$ and the initial point is $(x_{0},y_{0})=(0.5,0.625)$, for which the trajectory forms a torus.
Takens’s theorem [takens2006detecting] implies that, for a generic observable $f:M\to\mathbb{R}$, the topology of a smooth attractor can be recovered. For a trajectory $\{x(t)\}_{t\geq 0}$ in a system, the time-delay embedding or sliding window with delay $\tau>0$ and dimension $m\geq 1$ is
$SW(f;\tau,m)=\bigl\{\bigl(f(x(t)),f(x(t+\tau)),\dots,f(x(t+(m-1)\tau))\bigr)\in\mathbb{R}^{m}\mid t\in[0,T]\bigr\}.$
Following [charo2020topology], we apply this to the system (1) with the observable $f(x,y)=x$, delay $\tau=5$, and $m=4$, and obtain a finite subset of $\mathbb{R}^{4}$. The left part of Figure 4 illustrates the projection of the subset onto its first three coordinates.
We randomly select 1000 points from the subset and compute two persistence diagrams, one using the Euclidean distance and the other using the distance $d_{c}$ in Definition 1.1. The middle and right parts of Figure 4 display these diagrams. For the Euclidean case the shorter $1$-cocycle is not detected, whereas the diagram computed from the distance $d_{c}$ exhibits both one-dimensional homology classes and the unique two-dimensional homology class of the torus.
Figure 4: Projection of the time-delay embedding of the double-gyre system onto the first three coordinates (left); persistence diagrams for a subsample of time-delay embedded points with the Euclidean distance (middle) and with the distance $d_{c}$ (right).
4.2 Approximate quasi-halo orbit in the Saturn–Enceladus system
The circular restricted three-body problem (CR3BP) models the motion of a massless particle under the gravitational attraction of two massive bodies moving on circular orbits [szebehely2012theory, frauenfelder2018restricted]. In a rotating frame, the two primaries are fixed, and this system admits five equilibrium points, called Lagrange points. Near the collinear points $L_{1}$ and $L_{2}$, there exist non-planar periodic orbits called halo orbits [connor1984three].
For a dynamical system $(M,\phi)$ and the $k$-torus $\mathbb{T}^{k}$, a compact invariant submanifold $T\subset M$ is called a quasi-periodic torus of the flow $(\phi_{t})_{t\geq 0}$ if there exist a homeomorphism $h:T\to\mathbb{T}^{k}$ and a vector $\omega\in\mathbb{R}^{k}$ with rationally independent components such that $h(\phi_{t}(x))=h(x)+t\omega$ in $\mathbb{T}^{k}$ for every $t\geq 0$ and $x\in T$. We perturb a halo orbit near $L_{2}$ in the Saturn–Enceladus system, and obtain a Liouville torus that approximates the quasi-halo orbit.
For the Saturn–Enceladus system, we work in the phase space $\mathbb{R}^{6}=\{(q_{1},q_{2},q_{3},p_{1},p_{2},p_{3})\mid q_{i},p_{i}\in\mathbb{R},\,1\leq i\leq 3\}$. We approximate the quasi-halo orbit whose closure is diffeomorphic to the 3-torus $\mathbb{T}^{3}=S^{1}\times S^{1}\times S^{1}$ embedded in $\mathbb{R}^{6}$. Applying the surface-of-section method [poincare1893methodes, meiss2007differential], we collect the transverse intersections of this orbit with the section $\{q_{1}>L_{2},q_{2}=0\}\subset\mathbb{R}^{6}$. The support of the resulting point set is diffeomorphic to the 2-torus $\mathbb{T}^{2}=S^{1}\times S^{1}$ lying in $\mathbb{R}^{5}$.
We choose 4000 points from this point set. Since the central halo orbit of the quasi-periodic trajectory is close to degeneracy, the invariant torus exhibits significant distortions. Figure 5 shows the degree-0 and degree-$1$ persistence diagrams for this point cloud, computed using both the Euclidean distance and the distance $d_{c}$ in Definition 1.1. Because of the narrow bottleneck and large curvature of the torus, the persistence diagram based on the Euclidean distance contains only one prominent one-dimensional homology class, whereas the diagram based on $d_{c}$ exhibits both.
Figure 5: Three-dimensional projection of the points on the surface of section for the approximate quasi-halo orbit (left); degree-0 and degree-$1$ persistence diagrams for a subsample of these points, computed with the Euclidean distance (middle) and with the distance $d_{c}$ (right).
4.3 3D image shape
Topological data analysis is widely used in machine learning to characterize the geometric structure of datasets. Recent work integrates TDA into learning pipelines for classification, clustering, and anomaly detection [hensel2021survey, tauzin2021giotto]. We use the ModelNet40 dataset [wu20153d], which contains 12,311 CAD models across 40 categories. From this dataset we extract one finite subset from the car class and one from the vase class, which we display in the left part of Figure 6. Observe that the global structure of the point cloud of the car class appears closer to spherical than that of the vase class.
Figure 6: Visualizations of the 3D point clouds (left), and persistence diagrams of each dataset computed with the Euclidean distance (middle) and with the distance $d_{c}$ (right).
From each finite subset we sample 800 points to compute persistence diagrams with respect to both distances. To capture the coarse global structure, we only use these 800 points to estimate the tangent spaces, in contrast to the previous examples. The middle and right parts of Figure 6 display the resulting persistence diagrams. While the Euclidean distance does not separate the two datasets, the distance $d_{c}$ in Definition 1.1 reveals a nontrivial two-dimensional homology class for the car dataset.
Appendix A Proof of Theorem 3.2
Proof of Theorem 3.2.
Note that, by the Leibniz rule,
$\nabla_{v}^{M}(\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2})$ $=\nabla_{v}^{M}\langle\mathbf{II}(v,\cdot),\mathbf{II}(v,\cdot)\rangle_{\operatorname{HS}}$
$=2\langle\nabla^{M}_{v}(\mathbf{II})(v,\cdot),\mathbf{II}(v,\cdot)\rangle_{\operatorname{HS}}$
where we used $\nabla_{v}^{M}v=0$. On the other hand,
$\nabla_{v}^{M}(\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2})=2\cdot\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}\cdot\nabla_{v}^{M}(\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}).$
Therefore,
$\nabla_{v}^{M}(\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}})=\frac{\langle\nabla_{v}^{M}(\mathbf{II})(v,\cdot),\mathbf{II}(v,\cdot)\rangle_{\operatorname{HS}}}{\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}}.$
Hence
$\nabla_{v}^{M}(\log\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}})$ $=\frac{\nabla_{v}^{M}(\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}})}{\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}}$
$=\frac{\langle\nabla_{v}^{M}(\mathbf{II})(v,\cdot),\mathbf{II}(v,\cdot)\rangle_{\operatorname{HS}}}{\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}$
and applying the Cauchy–Schwarz inequality gives the desired result. ∎
Appendix B Details of Example 1
We keep the setting of Example 1 with $T^{2}\subset\mathbb{R}^{3}$ parametrized by
$(x,y,z)=\bigl((R+r\cos v)\cos u,\ (R+r\cos v)\sin u,\ r\sin v\bigr),\qquad u,v\in[0,2\pi),\ 0<r<R.$
We assume that $\frac{r}{R}\leq\frac{1}{m}$ for some $m\geq 2$, and set $c=(R-r)^{2}.$ Consider the loops
$\gamma_{1}=\{(R\cos u,R\sin u,r):u\in[0,2\pi]\},\quad\gamma_{2}=\{(R\cos u,R\sin u,-r):u\in[0,2\pi]\},$
$\gamma_{3}=\{(R+r\cos v,0,r\sin v):v\in[0,2\pi]\}.$
Under the identification $\pi_{1}(T^{2})\cong\mathbb{Z}^{2}$, we take $\gamma_{1}]=(1,0)$ and $\gamma_{3}]=(0,1)$. The shortest Euclidean representative of $\gamma_{1}]$ is
$\gamma_{4}=\{((R-r)\cos u,(R-r)\sin u,0):u\in[0,2\pi]\},$
whose Euclidean length is $2\pi(R-r)$. Hence any loop $\gamma\subset T^{2}$ with $\gamma]=(a,b)$ and $|a|\geq 1$ satisfies
$\mathrm{len}_{c}(\gamma)\geq\mathrm{len}(\gamma)\geq 2\pi(R-r).$
Next, we consider loops $\delta\subset T^{2}$ with $\delta]=(0,b)$ and $|b|\geq 1$. Such a loop meets both $\gamma_{1}$ and $\gamma_{2}$. Along $\delta$, the Gauss map $\bar{\mathbf{g}}^{+}$ traverses an arc in $S^{2}$ joining the antipodal points $\bar{\mathbf{g}}^{+}(\gamma_{1})$ and $\bar{\mathbf{g}}^{+}(\gamma_{2})$, so the Grassmannian component of $\delta$ has length at least $2\pi$. For the product metric, this implies
$\mathrm{len}_{c}(\delta)\geq 2\pi\sqrt{c}.$
Combining these two cases, we obtain $\mathrm{sys}_{c}(T^{2})\geq\min\bigl(2\pi(R-r),\,2\pi\sqrt{c}\bigr)$. With our choice $c=(R-r)^{2}$ we have $\sqrt{c}=R-r$, hence
$\mathrm{sys}_{c}(T^{2})\geq 2\pi(R-r).$
In the coordinates $(u,v)$ the first and second fundamental forms of $T^{2}\subset\mathbb{R}^{3}$ are
$g]=\begin{pmatrix}(R+r\cos v)^{2}&0\\[2.0pt] 0&r^{2}\end{pmatrix},\qquad[\mathbf{II}]=\begin{pmatrix}-(R+r\cos v)\cos v&0\\[2.0pt] 0&-r\end{pmatrix}$,
so
$g_{c}]=\begin{pmatrix}(R+r\cos v)^{2}+c\cos^{2}v&0\\[2.0pt] 0&r^{2}+c\end{pmatrix}$.
It follows that
$\mathrm{vol}_{c}(T^{2})=2\pi\sqrt{r^{2}+c}\int_{0}^{2\pi}\sqrt{(R+r\cos v)^{2}+c\cos^{2}v}\,dv.$
Using the inequality $\sqrt{A^{2}+B^{2}}\leq|A|+|B|$ with $A=R+r\cos v$ and $B=\sqrt{c}\cos v$, we obtain
$\mathrm{vol}_{c}(T^{2})\leq 2\pi\sqrt{r^{2}+c}\int_{0}^{2\pi}\bigl((R+r\cos v)+\sqrt{c}\,|\cos v|\bigr)\,dv=2\pi\sqrt{r^{2}+c}\,(2\pi R+4\sqrt{c}).$
Let $t=\frac{r}{R}\in(0,1)$. Then $c=(R-r)^{2}=R^{2}(1-t)^{2}$, and
$r^{2}+c=R^{2}\bigl(t^{2}+(1-t)^{2}\bigr),\qquad\sqrt{c}=R(1-t).$
Substituting these into the previous bounds gives
$\mathrm{sys}_{c}(T^{2})^{2}\geq 4\pi^{2}R^{2}(1-t)^{2},\qquad\mathrm{vol}_{c}(T^{2})\leq 2\pi R^{2}\sqrt{t^{2}+(1-t)^{2}}\,\bigl(2\pi+4(1-t)\bigr).$
Therefore
$\frac{\mathrm{sys}_{c}(T^{2})^{2}}{\mathrm{vol}_{c}(T^{2})}\geq\frac{2\pi(1-t)^{2}}{\sqrt{t^{2}+(1-t)^{2}}\,\bigl(2\pi+4(1-t)\bigr)}.$
Let $\nu$ be the unit normal vector field on $T^{2}$. For any bottleneck $(q_{1},q_{2})$ of $T^{2}\subset\mathbb{R}^{3}$,
$\|q_{1}-q_{2}\|_{\mathbb{R}^{3}}\geq\begin{cases}2r,&\nu_{q_{1}}=-\nu_{q_{2}},\\[2.0pt] 2R,&\nu_{q_{1}}=\nu_{q_{2}},\end{cases}$
and both bounds are attained. In the first case, the Gauss images are antipodal, and in the second they coincide, so
$L_{c}(T^{2})=\min\!\left(\tfrac{1}{2}\sqrt{\,4r^{2}+c\pi^{2}\,},\,R\right).$
With $c=(R-r)^{2}$ and $t=r/R$ this equality becomes
$L_{c}(T^{2})=\min\!\left(\tfrac{1}{2}\sqrt{\,4R^{2}t^{2}+R^{2}(1-t)^{2}\pi^{2}\,},\,R\right)$
and we deduce $L_{c}(T^{2})=R$ for sufficiently small $t>0$. Combining this with $\mathrm{vol}_{c}(T^{2})\leq 2\pi R^{2}\sqrt{t^{2}+(1-t)^{2}}\,\bigl(2\pi+4(1-t)\bigr)$ yields
$\frac{L_{c}(T^{2})^{2}}{\mathrm{vol}_{c}(T^{2})}\geq\frac{1}{2\pi\sqrt{t^{2}+(1-t)^{2}}\,\bigl(2\pi+4(1-t)\bigr)}.$
Appendix C Proof of Theorem 4.2
Before proving Theorem 4.2, we recall the following two theorems. Let $M\subset\mathbb{R}^{D}$ be a smooth closed connected orientable $d$-dimensional submanifold, and let $\mathsf{rch}_{\mathbb{R}^{D}}(M)=\tau>0$.
Theorem C.1 (Proposition 6.2 in [niyogi2008finding]).
Let $p,q\in M$, and let $\theta_{1}\geq\dots\geq\theta_{d}\geq 0$ be the principal angles between $T_{p}M$ and $T_{q}M$. Then
$\cos\theta_{1}\geq 1-\frac{1}{\tau}d_{M}(p,q).$
Theorem C.2 (Proposition 6.3 in [niyogi2008finding]).
If $p,q\in M$ satisfy $\|p-q\|_{\mathbb{R}^{D}}\leq\frac{\tau}{2}$, then
$d_{M}(p,q)\leq\tau-\tau\sqrt{1-\frac{2\|p-q\|_{\mathbb{R}^{D}}}{\tau}}.$
Proof of Theorem 4.2.
We claim that if $\|p-q\|_{\mathbb{R}^{D}}<\frac{\tau}{2}$, then no principal angle between $T_{p}M$ and $T_{q}M$ equals $\frac{\pi}{2}$. Assuming this claim, let $\gamma:[0,t]\mapsto M$ be a unit-speed geodesic from $p$ to $q$. Denote the principal angles between $T_{p}M$ and $T_{\gamma(s)}M$ as $\theta_{1}(s),\dots,\theta_{d}(s)$. Then
$\det(B_{p}^{\top}B_{\gamma(s)})=\pm\cos^{2}\theta_{1}(s)\cdot\cos^{2}\theta_{2}(s)\cdot\ldots\cdot\cos^{2}\theta_{d}(s).$
Since the map $s\mapsto\det(B_{p}^{\top}B_{\gamma(s)})$ is continuous, $\theta_{i}(0)=0$ for every $1\leq i\leq d$ so that $\det(B_{p}^{\top}B_{p})=1$, and $\theta_{i}(s)$ never becomes $\frac{\pi}{2}$ for every $0\leq s\leq t$, we have $\det(B_{p}^{\top}B_{\gamma(t)})=\det(B_{p}^{\top}B_{q})>0$.
To show the claim, assume that $\|p-q\|_{\mathbb{R}^{D}}<\frac{\tau}{2}.$ By Theorem C.1, for principal angles $\theta_{1},\dots,\theta_{d}$ between $T_{p}M$ and $T_{q}M$, we have
$1-\frac{1}{\tau}d_{M}(p,q)\leq\cos\theta_{1}\leq\dots\leq\cos\theta_{d}$
so that $\cos\theta_{i}\geq 1-\frac{1}{\tau}d_{M}(p,q)$ for every $1\leq i\leq d.$ Combining Theorem C.2 with the above and using the assumption, we have
$\cos\theta_{i}\geq\sqrt{1-\frac{2\|p-q\|_{\mathbb{R}^{D}}}{\tau}}>0.$
This proves the claim and the theorem. ∎
Author Information.
Dongwoo Gang
Department of Mathematical Sciences, Seoul National University
Email: dongwoo.gang@snu.ac.kr