### v2

# Oriented Grassmannian Bundle, Normal Curvature Reduction, and Persistent Homology

Abstract
We consider a smooth closed orientable submanifold $M\subset\mathbb{R}^{D}$ with narrow cycles. We embed $M$ into a scaled oriented Grassmannian bundle via the Gauss map in order to enlarge the scale of these cycles. Under mild assumptions, we show that this embedding reduces the normal curvature of the embedded submanifold in directions where the original normal curvature is large. For smooth closed hypersurfaces, we further show that this construction increases the distance between antipodal points of narrow cycles for fixed volume.
We then obtain an explicit range of radii for which the ambient Čech complex on this Grassmannian bundle is homotopy equivalent to the embedded manifold, yielding lower bounds on the scales at which the Čech filtration recovers the homology of $M$. Since the distance induced by the embedding depends on both positions and oriented tangent spaces, we work with Whitney $C^{1}$ convergence of embeddings and prove that the associated Čech persistent homology is stable with respect to the interleaving distance. Finally, we describe a procedure for computing a distance matrix for a finite subset with respect to this embedding and illustrate the construction on several examples, including an approximate quasi-halo orbit in the Saturn–Enceladus system.
1 Introduction
Persistent homology computes barcodes of filtered chain complexes constructed from a metric space $X$ [zomorodian2004computing, oudot2015persistence]. For $r>0$, the Čech complex $\check{C}(X;r)$ is the nerve of the open cover of $r$-balls $\{B_{X}(x,r)\}_{x\in X}$, and the Vietoris–Rips complex $\mathrm{VR}(X;r)$ is the abstract simplicial complex whose simplices are the finite subsets $\sigma\subset X$ with $\operatorname{diam}(\sigma)<r$. If $M$ is a Riemannian manifold, then the Čech complex $\check{C}(M;r)$ is homotopy equivalent to $M$ for $r<r_{\operatorname{conv}}(M)$ by the nerve lemma, while the systole and the filling radius bound the range of $r$ for which the Vietoris–Rips complex $\mathrm{VR}(M;r)$ has the same first and top-dimensional homology groups as $M$ [virk20201, lim2024vietoris, balitskiy2025geometric].
The reach (or normal injectivity radius) $\mathsf{rch}_{\mathbb{R}^{D}}(M)$ of a smooth submanifold $M\subset\mathbb{R}^{D}$ is the supremum of $\tau>0$ such that every point $x\in\mathbb{R}^{D}$ with $d_{\mathbb{R}^{D}}(x,M)<\tau$ has a unique nearest point in $M$ [federer1959curvature, aamari2019estimating]. For smooth embedded submanifolds, this agrees with the supremum of radii for which the normal exponential map is injective. The reach of $M$ gives bounds on the scales $r$ where the Vietoris–Rips or Čech complex built from sufficiently dense points of $M$ in the ambient Euclidean metric [niyogi2008finding, kim2019homotopy], and the metric thickening of $M$ [adams2019metric, adams2022metric, adams2024persistent], the 1-Wasserstein analogue of the Vietoris–Rips complex, are homotopy equivalent to $M$.
Aamari et al. [aamari2019estimating] showed that the reach of a submanifold $M\subset\mathbb{R}^{D}$ can be expressed as the minimum of the reciprocal of the maximal normal curvature of $M$ and the minimal width of its bottlenecks, pairs of points in $M$ whose connecting line segment is orthogonal to both tangent spaces. For example, for a highly eccentric ellipse in $\mathbb{R}^{2}$ or a thin $2$-torus with small systole in $\mathbb{R}^{3}$ as in Figure 1, the normal exponential map ceases to be injective at a radius comparable to the reach, since the points of large curvature (marked in red) induce focal points and the bottlenecks (marked in blue) lie close to each other in the Euclidean metric.
In both examples, the indicated bottlenecks have opposite normal directions. This suggests that, for an orientable submanifold $M\subset\mathbb{R}^{D}$, embedding $M$ into an augmented ambient space, such as the oriented Grassmannian bundle $\mathbb{R}^{D}\times\mathbf{Gr}^{+}(D,d)$, can enlarge the normal injectivity radius of $M$. In the product metric on $\mathbb{R}^{D}\times\mathbf{Gr}^{+}(D,d)$, the contribution from the Grassmannian factor enlarges the lengths of curves near points of large normal curvature, so that unit-speed geodesics have smaller normal curvature, and bottleneck pairs with opposite normal directions are pushed farther apart because their oriented tangent spaces are far from each other in $\mathbf{Gr}^{+}(D,d)$.
Figure 1: A highly eccentric ellipse (left) and a thin 2-torus (right). Points of large curvature are marked in red, and the narrowest bottlenecks are indicated in blue.
To formalize the above discussion, let $\iota:M\to\mathbb{R}^{D}$ be an embedding of a smooth closed orientable $d$-dimensional submanifold. Fix $c>0$ and consider the scaled oriented Grassmannian $(\mathbf{Gr}_{c}^{+}(D,d),g^{\mathbf{Gr}_{c}^{+}(D,d)})$, whose underlying manifold is the oriented Grassmannian $\mathbf{Gr}^{+}(D,d)$ equipped with the rescaled metric $g^{\mathbf{Gr}_{c}^{+}(D,d)}=c\cdot g^{\mathbf{Gr}^{+}(D,d)}$. Let $\bar{\mathbf{g}}^{+}$ denote the Gauss map, and define the rescaled Gauss map
$\bar{\mathbf{g}}^{+}_{c}:M\xrightarrow{\bar{\mathbf{g}}^{+}}(\mathbf{Gr}^{+}(D,d),g^{\mathbf{Gr}^{+}(D,d)})\xrightarrow{\cdot c}(\mathbf{Gr}_{c}^{+}(D,d),g^{\mathbf{Gr}_{c}^{+}(D,d)}).$
We consider the graph embedding $\iota\times\bar{\mathbf{g}}_{c}^{+}:M\to(\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d),\,g^{\mathbb{R}^{D}}\oplus g^{\mathbf{Gr}_{c}^{+}(D,d)}),$ and define $g_{c}=(\iota\times\bar{\mathbf{g}}^{+}_{c})^{*}(g^{\mathbb{R}^{D}}\oplus g^{\mathbf{Gr}_{c}^{+}(D,d)})$. Throughout, we write $\mathbf{II}$ for the second fundamental form of $M\subset\mathbb{R}^{D}$ with the induced Euclidean metric and $\mathbf{II}_{c}$ for that of $(M,g_{c})\subset\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$ with the product metric. Likewise, we write $\mathrm{vol}_{c}(M)$ for the Riemannian volume of $(M,g_{c})$ and $\mathrm{vol}(M)$ for the Euclidean volume of $M\subset\mathbb{R}^{D}$.
Definition 1.1.
For $c>0$, we define a distance $d_{c}$ on $M$ by
$d_{c}(q_{1},q_{2})=d_{\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)}\bigl((q_{1},\bar{\mathbf{g}}_{c}^{+}(q_{1})),(q_{2},\bar{\mathbf{g}}_{c}^{+}(q_{2}))\bigr)=\sqrt{\|q_{1}-q_{2}\|_{\mathbb{R}^{D}}^{2}+c\,d_{\mathbf{Gr}^{+}(D,d)}\bigl(\bar{\mathbf{g}}^{+}(q_{1}),\bar{\mathbf{g}}^{+}(q_{2})\bigr)^{2}},$
where $\|\cdot\|_{\mathbb{R}^{D}}$ denotes the Euclidean norm and $d_{\mathbf{Gr}^{+}(D,d)}$ is the Riemannian distance on the oriented Grassmannian $\mathbf{Gr}^{+}(D,d)$.
We first show that, by embedding $M$ into $\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$, we can reduce the normal curvature of $M$ along directions where the original normal curvature is large. For $p\in M$ and a nonzero vector $v\in T_{p}M$, denote by $\kappa_{c}(v)$ and $\kappa(v)$ the (operator norms of the) normal curvatures of $(M,g_{c})\subset\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d)$ at $(p,\bar{\mathbf{g}}_{c}^{+}(p))$ in the direction $(v,d\bar{\mathbf{g}}_{c}^{+}(v))$ and of $M\subset\mathbb{R}^{D}$ at $p$ in the direction $v$, respectively, where both geodesics are parametrized by arc length.
Theorem 1.1.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable $d$-dimensional submanifold. Fix $p\in M$ and a unit vector $v\in T_{p}M$. Assume that the normal curvature at $p$ in the direction $v$ is sufficiently large so that
$\mathrm{vol}_{c}(M)\leq\bigl(1+c\,\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}\bigr)^{d/2}\mathrm{vol}(M),\qquad\kappa(v)>\frac{\|(\nabla_{v}^{M}\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}}{\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}},$
where $\|\cdot\|_{\operatorname{HS}}$ denotes the Hilbert–Schmidt norm. Then, for every $c>0$,
$\kappa_{c}(v)\,\mathrm{vol}_{c}(M)^{1/d}\;\leq\;\frac{\sqrt{\kappa(v)^{2}+c\|(\nabla_{v}^{M}\mathbf{II})(v,\cdot)\|_{\operatorname{HS}}^{2}}}{1+c\|\mathbf{II}(v,\cdot)\|_{\operatorname{HS}}^{2}}\,\mathrm{vol}_{c}(M)^{1/d}\;<\;\kappa(v)\,\mathrm{vol}(M)^{1/d},$
and the middle term in the inequalities is strictly decreasing as $c$ grows.
In directions with large normal curvatures, even a small displacement along the curve already produces a large change in the tangent space, so the Grassmannian component of $g_{c}$ contributes an increasingly large portion of the speed. Since the normal curvature is defined as the size of the normal acceleration divided by the square of this speed, reparameterizing the curve to have unit speed with respect to $g_{c}$ makes the normal curvature decrease as $c$ grows.
Next, to quantify how distances between bottleneck pairs change with respect to $d_{c}$, we set
$L_{c}(M)$ $=\min\Bigl\{\frac{1}{2}d_{c}(q_{1},q_{2})\;\Big|\;(q_{1},q_{2})\text{ is a bottleneck of }M\subset\mathbb{R}^{D}\Bigr\}$
$=\min\Bigl\{\frac{1}{2}d_{c}(q_{1},q_{2})\;\Big|\;\overline{q_{1}q_{2}}\perp T_{q_{1}}M,\ \overline{q_{1}q_{2}}\perp T_{q_{2}}M\Bigr\}.$
Similarly, in the Euclidean case, we define
$L(M)=\min\Bigl\{\frac{1}{2}\|q_{1}-q_{2}\|_{\mathbb{R}^{D}}\;\Big|\;(q_{1},q_{2})\text{ is a bottleneck of }M\subset\mathbb{R}^{D}\Bigr\}.$
Under a mild curvature hypothesis, we show that for a smooth closed orientable hypersurface $M$, the bottlenecks are separated when distances are measured with $d_{c}$ from Definition 1.1.
Theorem 1.2.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable hypersurface. Suppose that $L(M)\leq\frac{1}{\|\mathbf{II}\|_{2}}$. Then for every $c\in\bigl(0,\tfrac{12L(M)^{2}}{\pi^{2}}\bigr]$,
$\frac{L_{c}(M)}{\mathrm{vol}_{c}(M)^{1/(D-1)}}\;\geq\;\frac{\sqrt{4L(M)^{2}+c\pi^{2}}}{2\,\mathrm{vol}_{c}(M)^{1/(D-1)}}\;>\;\frac{L(M)}{\mathrm{vol}(M)^{1/(D-1)}},$
and the middle term in the inequalities is strictly increasing in $c$ on the interval $\bigl(0,\tfrac{12L(M)^{2}}{\pi^{2}}\bigr]$.
We obtain an explicit range of radii $r$ for which the ambient Čech complex $\check{C}(M,\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d);r)$ is homotopy equivalent to $M$. In particular, this yields a lower bound on the lengths of the barcodes in the associated persistent homology, from which the homology of $M$ is recovered from the distance $d_{c}$.
Theorem 1.3.
Let $M\subset\mathbb{R}^{D}$ be a smooth closed orientable $d$-dimensional submanifold. For $c>0$, the ambient Čech complex $\check{C}(M,\mathbb{R}^{D}\times\mathbf{Gr}_{c}^{+}(D,d);r)$ is homotopy equivalent to $M$ for all
$r<\min\!\left(\sqrt{\frac{c}{2}}\arctan\sqrt{\frac{2}{c\|\mathbf{II}_{c}\|_{2}}},\,\frac{\sqrt{c}\,\pi}{2},\,L^{\prime}_{c}(M)\right),$
where
$L^{\prime}_{c}(M)=\min\Bigl\{\frac{1}{2}d_{c}(q_{1},q_{2})\;\Bigm|\;\overline{(q_{1},\bar{\mathbf{g}}_{c}^{+}(q_{1}))(q_{2},\bar{\mathbf{g}}_{c}^{+}(q_{2}))}\perp T_{(q_{i},\bar{\mathbf{g}}_{c}^{+}(q_{i}))}(M,g_{c}),\ i=1,2\Bigr\}.$
[…]
Theorem 1.4.
Let $\{M_{i}\}_{i\in\mathbb{N}}$ be a sequence of smooth closed orientable $d$-dimensional submanifolds of $\mathbb{R}^{D}$ converging to a smooth closed orientable submanifold $M_{\infty}\subset\mathbb{R}^{D}$ in the Whitney $C^{1}$ topology (see Definition LABEL:Whitney). Fix $c>0$. Then for every field $\mathbb{k}$ and every $j\in\mathbb{N}\cup\{0\}$,
[…]
where $d_{I}$ denotes the interleaving distance between persistence modules.
Based on the above theorems, we describe a method for computing the distance matrix for a finite subset of $M$ with respect to the distance $d_{c}$ from Definition 1.1. To illustrate the performance of this method, we display three computational examples: a time-delay embedded attractor, an approximate quasi-halo orbit in the Saturn–Enceladus system, and a classification of three-dimensional image shapes.
2 Theoretical background
2.1 Persistence theory
A filtered space (or filtration) $\mathbb{X}$ is a collection of topological spaces $\{X_{i}\}_{i\in\mathbb{R}_{\geq 0}}$ equipped with inclusion maps $\iota_{i}^{j}:X_{i}\to X_{j}$ for all $i\leq j$. Let $(X,d)$ be a compact metric space and let $\varepsilon\in\mathbb{R}_{\geq 0}$. The Vietoris–Rips complex $\mathrm{VR}(X;\varepsilon)$ is the abstract simplicial complex whose simplices are finite subsets $\sigma\subset X$ such that $d(x,y)<\varepsilon$ for all $x,y\in\sigma$. For a non-decreasing sequence of parameters $\varepsilon_{0}\leq\varepsilon_{1}\leq\varepsilon_{2}\leq\dots$, these complexes form a filtered space
$\mathrm{VR}(X;\varepsilon_{0})\hookrightarrow\mathrm{VR}(X;\varepsilon_{1})\hookrightarrow\mathrm{VR}(X;\varepsilon_{2})\hookrightarrow\cdots,$
[…]
For $n\in\mathbb{N}\cup\{0\}$, a field $\mathbb{k}$ and a filtration $\mathbb{X}=\{X_{i}\}_{i\in\mathbb{R}_{\geq 0}}$, the $n$-th persistent homology $H_{n}(\mathbb{X};\mathbb{k})$ of $\mathbb{X}$ is a family of vector spaces $H_{n}(X_{i};\mathbb{k})$ with induced linear maps $H_{n}(\iota_{i}^{j}):H_{n}(X_{i};\mathbb{k})\to H_{n}(X_{j};\mathbb{k})$ for $i\leq j$. If $H_{n}(\mathbb{X};\mathbb{k})$ is pointwise finite-dimensional and decomposes as a direct sum of interval modules, then it corresponds to a multiset of points $(b,d)\subset\mathbb{R}_{\geq 0}^{2}$, called the persistence diagram in degree $n$.
For a field $\mathbb{k}$, a persistence module (over $\mathbb{R}_{\geq 0}$) $\mathbb{V}$ is a family of vector spaces $\{V_{i}\}_{i\in\mathbb{R}_{\geq 0}}$ equipped with linear maps $v_{i}^{j}:V_{i}\to V_{j}$ for every $0\leq i\leq j$, satisfying $v_{j}^{k}\circ v_{i}^{j}=v_{i}^{k}$ whenever $i\leq j\leq k$ and $v_{i}^{i}$ is the identity map on $V_{i}$. A morphism of degree $\varepsilon$ between two persistence modules $\mathbb{V}$ and $\mathbb{W}$ is a family of linear maps $\Phi=\{\phi_{i}:V_{i}\to W_{i+\varepsilon}\}_{i\in\mathbb{R}_{\geq 0}}$ such that for all $i\leq j$, it holds that
$w_{i+\varepsilon}^{j+\varepsilon}\circ\phi_{i}\;=\;\phi_{j}\circ v_{i}^{j}.$
For $\varepsilon\geq 0$, the shift $\Sigma^{\varepsilon}$ of $\mathbb{V}$ by $\varepsilon$ is
$(\Sigma^{\varepsilon}\mathbb{V})_{i}=V_{i+\varepsilon},\qquad(\Sigma^{\varepsilon})(v_{i}^{j})=v_{i+\varepsilon}^{j+\varepsilon}:V_{i+\varepsilon}\to V_{j+\varepsilon}.$
Two persistence modules $\mathbb{V}$ and $\mathbb{W}$ are $\varepsilon$-interleaved if there exist morphisms $\Phi=\{\phi_{i}:V_{i}\to W_{i+\varepsilon}\}_{i\in\mathbb{R}_{\geq 0}}$ and